### v1

# Exploiting Negative Capacitance for Unconventional Coulomb Engineering

Abstract
It is known that the many-body ground state of a two-dimensional electron system can be tuned through Coulomb engineering by controlling the permittivity of the surrounding media. However, permittivities are traditionally restricted to positive values. In this paper we argue that the negative capacitance effect demonstrated in appropriately engineered structures can open new vistas in Coulomb engineering. Negative permittivities transform the natural repulsive interaction of electrons into an attractive one raising the intriguing possibility of a superconducting ground state. Using models of two-dimensional electron systems with linear and parabolic dispersion relations coupled to environments with negative capacitance, we estimate the strength and sign of the engineered Coulomb interaction and outline parameter regimes that could stabilize correlated electronic phases.
preprint: APS/123-QED
I Introduction
Coulomb interactions between electrons govern a variety of many-body phenomena, from stabilizing interacting phases to shaping the electrical and optical responses of matter. Consequently, several approaches have been developed to highlight the effects driven by Coulomb interactions, bringing them to the forefront of condensed matter physics. For example, one prominent approach [bistritzer2011moire] focuses on revealing Coulomb interactions by suppressing competing kinetic energies in materials engineered to host flat energy bands. This flat-band engineering technique has successfully stabilized interacting phases ranging from superconductivity [cao2018, lu2019superconductors, balents2020superconductivity, oh2021] to the fractional quantum anomalous Hall state [sharpe2019qahe, serlin2020qahe, li2021qahe] in two-dimensional materials [kennes2021moirereview]. However, because this approach does not directly modify Coulomb interactions, the resulting interacting phases are typically stabilized only at low temperatures.
A complementary strategy—known as Coulomb engineering—enables the direct tuning of Coulomb interactions by designing heterostructures with tailored electromagnetic environments [raja2017, song2018]. Although Coulomb engineering has successfully altered the electrical and optical responses of materials [jang2008graph, hwang2012fvr, yu2013qcap, kajino2019coulomb, whelan2020fvr], it remains challenging to tune Coulomb interactions over a wide range and to design interaction-driven phases of matter within this framework. This limitation arises because conventional electromagnetic environments explored so far are largely constrained by the requirement that their static dielectric function (permittivity) must remain positive and be greater than that of vacuum.
This work is motivated by recent experimental demonstrations of Negative Capacitance (NC) [zubko2016, yadav2019, das2021local]. We propose to modify the structures in those works to embed two-dimensional electron systems (2DES) such that the normally repulsive interaction between electrons is rendered attractive, thereby unveiling a new regime of unconventional Coulomb engineering.
The space of applications unlocked by this extended tuning of Coulomb interactions spans both quantitative modifications and qualitative transformations. On the quantitative side, one can expect changes to transport properties and optical response in the normal state of 2DES, potentially exceeding limits of conventional Coulomb engineering [jang2008graph, hwang2012fvr, yu2013qcap, raja2017, song2018, kajino2019coulomb, whelan2020fvr]. On the qualitative side, entirely new phases of matter may emerge [van2023cemott], especially once the customarily repulsive electron–electron interactions are rendered attractive. Of particular interest is superconductivity, which requires the formation of paired electronic states [bcs1957]. We provide estimates of pairing strength through the dimensionless parameter $\lambda-\mu^{*}$, adapted from electron–phonon theory [tolmachev1961, morelanderson, colemanbook], and find that values $\gtrsim 0.1$ are achievable in realistic regimes. Crucially, this parameter is highly tunable within our proposal, suggesting that NC-enabled Coulomb engineering could provide a new route toward stabilizing superconductivity and other ordered electronic phases, potentially even at elevated temperatures.
II Central Idea
To illustrate the main concept, we begin with the idealized structure shown schematically in Fig. 1(a) (connections to real material platforms are made in later sections). A two-dimensional electron system (2DES) is sandwiched between a gated conventional dielectric (DE) of thickness $L_{d}$ and a gated NC material of thickness $L_{nc}$. We model the DE and NC regions with diagonal permittivity tensors $\varepsilon^{nc(d)}=\text{diag}(\varepsilon^{nc(d)}_{\perp},\varepsilon^{nc(d)}_{\perp},\varepsilon^{nc(d)}_{z})$, where $z$ denotes the stacking direction and the superscripts $nc$ and $d$ represent NC and DE regions, respectively.
In the 2DES plane, the static Coulomb interaction $V_{\text{eff}}$ as a function of in-plane wavenumber $q$ is given by
$V_{\text{eff}}(q)=\frac{e^{2}}{2\varepsilon_{0}q}\cdot\frac{1}{\varepsilon_{\text{eff}}(q)}=\frac{v_{q}}{\varepsilon_{\text{eff}}(q)},$ (1)
where the effective dielectric function $\varepsilon_{\text{eff}}(q)$ is obtained self-consistently as
[…]
  • The expression for $\varepsilon_{\text{env}}(q)$ reflects the “anisotropic gate screening” effect, derived from Poisson’s equation (see Appendix A).
  • For the electron polarization function $\Pi_{el}$, we use the static limit of the Lindhard function for a 2DES [giuliani2008quantum], which reduces to the Thomas-Fermi (TF) form for wavenumber $q<2k_{F}$—the range relevant for this work.
Figure 1: Central idea. (a) System schematic - a two-dimensional electron system (2DES) surrounded by dielectric (conventional) and negative capacitance (NC) media, with respective thicknesses $L_{d}$ and $L_{nc}$. When the NC material is a ferroelectric, this configuration will be referred to as a Metal-Ferroelectric-2DES-Insulator-Metal (MF2IM) structure. (b) Tunability of the engineered 2DES Coulomb interaction $V_{\text{eff}}$ in the long-wavelength approximation as per Eq. (3), shown as a function of geometric capacitances $C_{nc}$ and $C_{d}$, and quantum capacitance $C_{q}$. The dashed black line indicates the limit of conventional Coulomb engineering ($C_{nc}>0$, region I). Notably, $V_{\text{eff}}$ can become negative (region III) if $C_{nc}$ is sufficiently negative.
Long-wavelength limit: Introducing dimensionless anisotropy factors $\eta_{nc}=\sqrt{\varepsilon^{nc}_{\perp}/\varepsilon^{nc}_{z}}$ and $\eta_{d}=\sqrt{\varepsilon^{d}_{\perp}/\varepsilon^{d}_{z}}$, we illustrate the central idea of NC-enabled unconventional Coulomb engineering by first considering the limit where $q\eta_{nc}L_{nc}\ll 1$ and $q\eta_{d}L_{d}\ll 1$. In this regime, the wavelengths associated with low-energy scattering of electrons in the 2DES are large compared to the gate distances. The assumption of $q$-independent permittivities in this section additionally requires that this wavelength is large compared to characteristic length scales of the polarizable DE and NC media (elaborated in later sections). Under these conditions, Eq. (2) reduces to a simple form:
$V_{\text{eff}}\cong\frac{e^{2}}{C_{nc}+C_{d}+C_{q}},$ (3)
where $C_{nc}$ and $C_{d}$ are geometric capacitances per unit area, and $C_{q}$ is the quantum capacitance of the 2DES:
$C_{nc}=\frac{\varepsilon_{0}\varepsilon^{nc}_{z}}{L_{nc}}\hskip 5.69046pt;\hskip 5.69046ptC_{d}=\frac{\varepsilon_{0}\varepsilon^{d}_{z}}{L_{d}}\hskip 5.69046pt;\hskip 5.69046ptC_{q}=2\varepsilon_{0}k_{\text{TF}}.$ (4)
Throughout this work, we set the ‘background’ dielectric constant of 2DES to unity for simplicity.
Eq. (3) encapsulates our central thesis: the inclusion of $C_{nc}<0$ enables the design of structures where $V_{\text{eff}}$ can be negative and engineered to have large magnitude values through careful matching of the constituent capacitances.
Before discussing the scope of this result, we first present a stability condition which constrains the values $C_{nc}$ can assume in the proposed configuration.
Structural Stability: The NC state of a material in isolation is thermodynamically unstable, but as argued in Ref. [salahuddindatta2008], it can be locally stabilized in a constituent layer provided the composite structure is stable against charge fluctuations.
We refer to the structure in Fig. 1(a) as Metal-Ferroelectric-2DES-Insulator-Metal (MF2IM) configuration (specifying NC material as ferroelectric to connect with notation used in literature [park2019modeling, luk2022ferroelectric]). The necessary and sufficient condition for stability of NC in the MF2IM configuration is (assuming $C_{nc}<0$ and $C_{d},C_{q}>0$, see Appendix B for details):
$C_{nc}+C_{d}+C_{q}<0.$ (5)
Setting $C_{q}=0$ recovers the familiar stability condition for the widely studied Metal-Ferroelectric-Insulator-Metal (MFIM) structure [salahuddindatta2008, park2019modeling, iniguez2019ferroelectric]. In the opposite limit of $C_{q}\to\infty$, the MF2IM structure decouples into two separate capacitors (assuming the 2DES is connected to a charge reservoir), and NC cannot be stabilized in this case.
Scope: Fig. 1(b) illustrates the scope of Eq. (3) by treating the capacitances as free variables. In conventional Coulomb engineering where all geometric capacitances are positive, the tunability of $V_{\text{eff}}$ is limited to region I (assuming $C_{q}>0$). The inclusion of $C_{nc}<0$ opens up new regimes, indicated as regions II and III. Most strikingly, in region III the sign of $V_{\text{\text{eff}}}$ can be reversed while still satisfying the stability condition in Eq. (5), transforming the naturally repulsive electron-electron interaction into an effective attraction.
Figure 2: Critical length scales. (a) Definition of two regimes based on three length scales. The wavelength of low-energy electron scattering in the 2DES relative to $2d$ (twice the domain width) in ferroelectric periodic domain texture (PDT) governs whether dispersion $\varepsilon^{f}_{z}(q)$ is important in extrinsic NC systems ($2d$ should be replaced by an appropriate lattice-related length scale in the case of intrinsic NC) and relative to $|\eta_{f}|L_{f}$ determines whether the linear approximation of gate screening effect is valid. (b), (c) Illustration of effective Coulomb interaction energy $V_{\text{eff}}$ as a function of wavenumber $q$ as per Eqs. (1) and (2), shown for linear and parabolic 2DES with typical $v_{F}$ and $m^{*}$ values. In both cases, 2DES carrier density $n$ was varied to access the long-wavelength (low $n=10^{15}/\text{m}^{2}$) and short-wavelength (high $n=10^{17}/\text{m}^{2}$) regimes. Note that changing $L_{f}$ (fixed at $4~\text{nm}$ for this figure) is an alternate way to access these regimes.
III Model details
In this section, we provide details of the models we use to estimate parameters appearing in Eq. (2), treating the 2DES carrier density $n$ and the NC material thickness $L_{nc}$ as tunable variables.
NC material: Within the growing literature on negative capacitance, it is useful to distinguish between transient and stabilized NC. Transient NC, such as that reported in [khan2015transient, khan2017transient, hoffmann2018transient, hoffmann2019unveiling, kim2019transient, qiao2021both], occurs during polarization switching and is therefore a non-equilibrium associated with dynamical relaxation of the order parameter. In contrast, stabilized NC [appleby2014stable, gao2014stable, zubko2016, hoffmann2018stabilization, yadav2019, das2021local, qiao2021both, qiao2024stable] refers to the quasi-static, near-equilibrium regime, where the system exhibits static negative permittivity. Beyond this dynamical classification, NC systems can also be categorized by material class— perovskite (such as [yadav2019]), oxide (such as [hoffmann2022intrinsic]), van der Waals (such as [wang2019vdw])—and by the underlying ferroic order, with both ferroelectrics [iniguez2019ferroelectric] and antiferroelectrics [qiao2021both, hoffmann2022antiferroelectric] being reported to exhibit NC behavior.
In the remainder of this work we restrict attention to stabilized NC, although the idea outlined in the previous section could be explored in the transient case as well. Within stabilized NC, two classes of systems are often distinguished: (i) extrinsic NC [zubko2016, yadav2019, das2021local], where the phenomenon is associated with multi-domain configurations and domain wall motion [junquera2023topological], and (ii) intrinsic NC [hoffmann2022intrinsic], where the effect appears to be independent of specific domain configurations. The theory developed in the previous section and summarized in Fig. 1 is formulated without assuming either classification, but we adopt a domain-provided model for extrinsic NC when performing calculations for specific material systems (results in Figs. 2 and 3).
We now provide details about the domain-provided model for stabilized NC in perovskite ferroelectric $\text{PbTiO}_{3}$. This choice is motivated by substantial evidence from experiments [zubko2016, yadav2019, das2021local] and corresponding theoretical explanation [luk2018prb, luk2019harnessing] providing a framework for modeling near-equilibrium NC. In ferroelectric $\text{PbTiO}_{3}$ ($\varepsilon^{nc}_{z(\perp)}\to\varepsilon^{f}_{z(\perp)}$, $L_{nc}\to L_{f}$), the polarization response is highly anisotropic, with $\varepsilon^{f}_{z}<0$ and $\varepsilon^{f}_{\perp}>0$. Starting from a periodic domain texture (PDT) as the equilibrium polarization configuration, the static negative permittivity $\varepsilon^{f}_{z}$ is explained as an overscreening effect due to the role of depolarization field in this system [luk2018prb, luk2019harnessing]. Assuming small deviations from equilibrium, the PDT dynamics can be described by an oscillator model for domain wall displacements, where $P$ is the electric polarization and $E_{\text{tot}}$ is the total electric field (external + depolarization):
$\ddot{P}(t)+\Gamma\dot{P}(t)+(\omega_{0}^{2}-\Omega^{2})P(t)=\Omega^{2}\cdot\varepsilon_{0}\varepsilon^{f}_{z,hf}E_{\text{tot}}(t).$ (6)
Here, $\Gamma$ is the damping constant for domain wall motion. The restoring force originates from Coulomb energy associated with excess surface charges when the PDT is displaced from equilibrium. Fourier analysis of the domain structure within Kittel model [kittel1946, junquera2023topological] yields the following expressions for frequency parameters [sidorkin2006domain, luk2018prb]:
$\begin{aligned} \omega_{0}^{2}=\frac{8P_{s}^{2}}{\pi\varepsilon_{0}\sqrt{\varepsilon^{f}_{\perp}\varepsilon^{f}_{z,\text{hf}}}ML_{f}}&\mathrm{ln}\Bigg(\mathrm{cosh}\Bigg(\sqrt{\frac{\varepsilon^{f}_{\perp}}{\varepsilon^{f}_{z,\text{hf}}}}\frac{\pi}{2}\frac{L_{f}}{d}\Bigg)\Bigg)\\ \Omega^{2}=&\dfrac{4P_{s}^{2}}{\varepsilon_{0}\varepsilon^{f}_{z,\text{hf}}Md}.\end{aligned}$ (7)
Here, $\omega_{0}$ is the characteristic oscillation frequency of the PDT system and $\Omega$ represents the strength of coupling to external electric field.
In a more detailed model, these frequency parameters should exhibit dispersion $\omega_{0}(q),\Omega(q)$ with additional indices for excitation type (longitudinal vs transverse) and branch (acoustic vs optic). The expressions in Eq. (7) account only for the longitudinal optic branch at zero wavenumber (see Appendix C), based on the expectation that this mode contributes dominantly to the NC effect in the long-wavelength limit. We refer to this as the ‘single-mode’ approximation.
Parameters for ferroelectric $\text{PbTiO}_{3}$ are chosen as follows [luk2018prb] : $P_{s}=0.65~\text{Cm}^{-2}$ is the spontaneous polarization, and $M$ is the domain wall mass per unit area estimated using interpolation formula $M=1.3\sqrt{L_{f}[\text{nm}]}\times 10^{-9}~\text{kgm}^{-2}$ [luk2018prb, zhangponomareva]. The equilibrium domain width $d$ is calculated using the Kittel model (explicit formula in Appendix C). $\varepsilon^{f}_{z,hf}=100$ is the “high-frequency” dielectric constant which captures the background polarizability of $\text{PbTiO}_{3}$ arising from other polarization mechanisms in the material, and $\varepsilon^{f}_{\perp}=30$.
The static linear response dielectric function corresponding to Eq. (6) is:
$\varepsilon^{f}_{z}\equiv\varepsilon^{f}_{z}(q=0,\omega=0)=\ \varepsilon^{f}_{z,\text{hf}}\cdot\frac{\omega_{0}^{2}}{\omega_{0}^{2}-\Omega^{2}}.$ (8)
Since $\omega_{0}<\Omega$ for ferroelectric $\text{PbTiO}_{3}$ in this model, the above expression yields a negative $\varepsilon^{f}_{z}$.
Dielectric: For the conventional dielectric part of environment (see Fig. 1(a)), we adopt values for hexagonal boron nitride (hBN) as a representative ‘low-k’ dielectric. Specifically, we set $\varepsilon^{d}_{z}=3.4$, $\varepsilon^{d}_{\perp}=6.86$ based on values reported in literature [ohbahbn, laturiahbn, Pierrethbn, kimhbn]. In Fig. 3(b), we choose hafnium oxide ($\text{HfO}_{2}$) as a representative ‘high-k’ dielectric in the long-wavelength regime with $\varepsilon^{d}_{z}=25$. The dielectric thickness is fixed at $L_{d}=4~\text{nm}$ in all cases.
2DES: The bare 2DES (in vacuum, $\varepsilon_{\text{env}}=1$) is modeled as a single-valley non-interacting electron gas, with either linear or parabolic low-energy dispersion. The two cases are characterized respectively by a Fermi velocity $v_{F}$, or an effective mass $m^{*}$. Spin and valley degeneracies are set to $g_{S}=2,g_{V}=1$ respectively. The quantum capacitance $C_{q}$ and Thomas-Fermi screening wavenumber $k_{\text{TF}}$ are
$\begin{aligned} \text{Linear:}\hskip 5.69046ptC_{q}=2\varepsilon_{0}k_{\text{TF}}&=\frac{e^{2}k_{F}}{\pi\hbar v_{F}},\\ \text{Parabolic:}\hskip 5.69046ptC_{q}=2\varepsilon_{0}k_{\text{TF}}&=\frac{e^{2}m^{*}}{\pi\hbar^{2}}.\end{aligned}$ (9)
Note that in the linear case $C_{q}$ and $k_{\text{TF}}$ scale with carrier density as $E_{F}$ ($\propto\sqrt{n}$). In the parabolic case, $C_{q}$ and $k_{TF}$ are independent of $n$.
IV Results
In this section, we apply the models described above to identify critical length and energy scales in the composite structure and estimate the strength of engineered Coulomb interactions in the low-energy regime.
Length scales: Let us first note that the quantity $\eta_{nc}\equiv\sqrt{\varepsilon^{nc}_{\perp}/\varepsilon^{nc}_{z}}$ is imaginary for an anisotropic NC material like $\text{PbTiO}_{3}$ which has $\varepsilon^{f}_{z}<0~,~\varepsilon^{f}_{\perp}>0$. Consequently the tanhtanh functions in Eq. (2) become tantan functions ($\tanh(ix)=i\tan(x)$):
$\varepsilon_{\text{env}}(q)=\frac{1}{2}\bigg[\frac{|\eta_{f}|\varepsilon^{f}_{z}}{\tan\big(q\hskip 0.85355pt|\eta_{f}|L_{f}\big.)}+\frac{\eta_{d}\varepsilon^{d}_{z}}{\tanh\big(q\hskip 1.42271pt\eta_{d}L_{d}\big.)}\bigg],$ (10)
where $|\eta_{f}|$ denotes the complex magnitude, and we have assumed that the dielectric part is conventional ($\varepsilon^{d}_{z},\varepsilon^{d}_{\perp}>0$). This form reveals that in the case of anisotropic NC the calculated $V_{\text{eff}}$ is periodic, and it is natural to identify the corresponding critical length scale $|\eta_{f}|L_{f}$. For a system with isotropic NC ($\varepsilon^{f}_{z}<0$, $\varepsilon^{f}_{\perp}<0$) $\eta_{f}$ is real-valued, the corresponding gate screening function remains tanhtanh, and $V_{\text{eff}}(q)$ is not periodic.
In addition to $|\eta_{f}|L_{f}$, two important length scales in the composite structure are $2k_{F}$ (low-energy scattering shell in the 2DES) and a characteristic length scale corresponding to the NC phenomenon, taken to be the domain width $d$ of PDT in the extrinsic case. We define two distinct regimes based on the relative values of these length scales: long-wavelength and short-wavelength.
In the long-wavelength limit, two simplifications emerge: wavelengths of electron scattering in 2DES are too large to resolve microscopic details on the scale of domain width $d$ (so that a simplified model with constant $\varepsilon^{nc}(q)$ suffices), and gate screening reduces to its linear limit since the gates see an effectively uniform charge distribution in the 2DES. Expressed as mathematical conditions:
  • In the long-wavelength regime, $2k_{F}d\ll 1$ and $2k_{F}|\eta_{f}|L_{f}\ll 1$, so that the single-mode approximation discussed in previous section is a reasonable starting point, and the simplified description of Eq. (3) is applicable [see Fig. 2(b)].
  • In the short-wavelength regime, $2k_{F}d\gg 1$ and $2k_{F}|\eta_{f}|L_{f}\gg 1$ , indicating that it is important to model dispersion $\varepsilon^{f}_{z}(q)$ (single-mode approximation may not necessarily be reliable) and gate screening modifies the $q$-space structure of $V_{\text{eff}}$ significantly [see Fig. 2(c)].
Figs. 2(b),(c) demonstrate how the two length scale regimes are accessed using 2DES carrier density $n$ as the control parameter in a model system with ferroelectric $\text{PbTiO}_{3}$ as the NC material. $V_{\text{eff}}$ as a function of wavenumber $q$ is shown. Ferroelectric thickness $L_{f}$ could be varied at fixed $n$ to achieve the same effect. The long-wavelength limit is accessed in the case of low carrier density and small thickness.
Energy scales: The low-energy picture described above is appropriate when the adiabatic condition $\hbar\omega_{0}\ll E_{F}$ is satisfied. Fig. 3(a) shows the ratio of 2DES Fermi energy $E_{F}$ to PDT oscillation energy $\hbar\omega_{0}$ (for $4~\text{nm}$ thick $\text{PbTiO}_{3}$) as a function of carrier density. At low density, the adiabatic condition begins to be satisfied ($E_{F}\sim 10\hbar\omega_{0}$) only for linear 2DES with high $v_{F}$. At high density, both linear and parabolic 2DES comfortably satisfy this condition ($E_{F}\sim 100\hbar\omega_{0}$) for typical $v_{F},m^{*}$ values.
The theory developed in this work is most applicable in the low-energy, long-wavelength regime. For this reason, we focus in the remainder of this section on the case of linear 2DES at low density, which according to Fig. 3(a) could be treated as a low-energy system for large enough $v_{F}$ and according to Fig. 2(b) could be modeled reasonably in the long-wavelength approximation (using Eq. (3)).
Figure 3: Critical energy scales and pairing strength. (a) Fermi energy $E_{F}$ of the 2DES, expressed in units of PDT single-mode energy $\hbar\omega_{0}$ for ferroelectric $\text{PbTiO}_{3}$, plotted as a function of Fermi wavenumber $k_{F}$ normalized to the ferroelectric thickness $L_{f}=4~\text{nm}$. The long-wavelength regime described in Fig. 2 is accessed in the limit of low density, shown as encircled region. (b) Pairing strength parameter $\lambda-\mu^{*}$ calculated for a linear 2DES in the long-wavelength approximation as a function of balancing parameter $\zeta$ defined in Eq. (13), shown for two different values of Fermi velocity $v_{F}$. Here, carrier density was fixed at $n=1\times 10^{15}/\text{m}^{2}$. Inset highlights where typical material systems (NC/DE) fall, according to the models used in this work.
From the perspective of real systems, the high density regime and case of parabolic 2DES are just as likely to exhibit interesting physics, but careful treatment of $V_{\text{eff}}(q,\omega)$ in the short-wavelength, high-energy regime is required and beyond the scope of this work.
Pairing strength: To quantify the strength of low-energy electron pairing in the 2DES, we use the parameters $\lambda$ and $\mu^{*}$, adapted from the theory of electron-phonon interaction and superconductivity [colemanbook]. To explain their definition in our context, we begin by decomposing the self-consistently screened interaction $V_{\text{eff}}$ in the following manner, based on the expected slowness of NC-mediated screening relative to direct screening by electrons ($\hbar\omega_{0}\ll E_{F}$):
$\begin{aligned} V_{\text{eff}}(q)&=v_{q}+v_{q}\Pi_{\text{el}}(q)V_{\text{eff}}(q)+v_{q}\Pi_{\text{env}}(q)V_{\text{eff}}(q)\\ &=\hskip 22.76228pt\underbrace{V_{\text{el}}(q)}_{\text{fast}}\hskip 22.76228pt+\underbrace{V_{\text{el}}(q)\Pi_{\text{env}}(q)V_{\text{eff}}(q)}_{\text{slow}},\end{aligned}$ (11)
where $\Pi_{\text{env}}(q)=\Pi_{nc}(q)+\Pi_{d}(q)$ and $V_{\text{el}}(q)=v_{q}/(1-v_{q}\Pi_{\text{el}}(q)$). We call the ‘slow’ term in Eq. (11) $\tilde{V}_{\text{env}}$ and calculate the following dimensionless constants [colemanbook]:
$\begin{aligned} \lambda=-N_{0}\langle\tilde{V}_{\text{env}}\rangle=&-N_{0}\times\frac{\int_{0}^{2\pi}d\theta~\tilde{V}_{\text{env}}(q(\theta))}{2\pi},\\ \mu=N_{0}\langle V_{el}\rangle&=N_{0}\times\frac{\int_{0}^{2\pi}d\theta~V_{el}(q(\theta))}{2\pi},\\ \mu^{*}=&~\frac{\mu}{1+\mu\ln(\frac{E_{F}}{\hbar\omega_{0}})},\end{aligned}$ (12)
where angle brackets denote averaging over Fermi surface (FS), $q(\theta)$ is appropriately defined for the 2D system, and $N_{0}$ is the density of states at $E_{F}$ per spin per unit area. The Coulomb pseudopotential $\mu^{*}$ [morelanderson] represents a renormalization of direct Coulomb repulsion ($\mu$) due to the large bandwidth for electron-electron scattering ($\sim E_{F}$) compared to environment-mediated screening ($\sim\hbar\omega_{0}$). Carrier pairing at accessible temperatures is expected when $\lambda-\mu^{*}>0$ and sufficiently large ($\gtrsim 0.1$).
In the long-wavelength regime, the problem can be fully specified by fixing values for the 2DES $n$ and $v_{F}$, and introducing a balancing parameter $\zeta$ defined as follows:
$C_{nc}+C_{d}=-\zeta C_{q}.$ (13)
To satisfy the long-wavelength stability condition in Eq. (5) we assume $\zeta>1$.
Fig. 3(b) shows calculated values of $\lambda-\mu^{*}$ for NC $\text{PbTiO}_{3}$ and linear 2DES ($n=1\times 10^{15}~/\text{m}^{2}$) in the long-wavelength regime. We emphasize two points from this figure. First, the values of $\lambda-\mu^{*}\sim 0.2-0.3$ shown in the inset, calculated using the models described in previous section, are sufficiently large to motivate experimental efforts to observe pairing in such systems. For comparison, in monolayer graphene the value of $\lambda$ arising from other pairing mechanisms such as the electron-phonon interaction is many orders of magnitude smaller [einenkel2011graphene]. Second, balancing the capacitances of NC, DE and 2DES offers a new route to engineer the value of $\lambda-\mu^{*}$. Particularly, as the balancing parameter approaches $\zeta=1$ from above, pairing can be pushed into stronger coupling regime.
Design considerations: From the perspective of NC stabilization, linear and parabolic 2DES systems present qualitative differences. For a linear 2DES, $C_{q}\propto\sqrt{n}/v_{F}$, so the choice of carrier density directly affects the condition for NC stabilization (Eq. (5)). For a parabolic 2DES, $C_{q}$ is independent of carrier density, and scales with effective mass as $C_{q}\propto m^{*}$. This means that effective mass is the critical parameter and that platforms with a larger $m^{*}$ parabolic 2DES will require larger negative $\varepsilon^{nc}$ for NC to be stabilized.
From the perspective of carrier pairing represented by $\lambda-\mu^{*}$, the functional dependence on $n$ and $L_{f}$ is more complicated due to the expressions for FS averaging and retardation effect ($\mu^{*}$). The essential route to maximizing interaction strength is to balance the positive and negative contributions in Eq. (2), and in the long-wavelength regime the problem simplifies considerably, leading to the points made in Fig. 3(b).
V Discussion
In summary, we have presented a framework for Coulomb engineering using negative capacitance materials. Key results include: the idea of Coulomb Engineering using NC (Fig. 1), a stability condition for NC in MF2IM configuration Eq. (5), categorization into regimes based on critical length scales (Fig. 2) and an estimation of energy scales and interaction strength for models of material systems (Fig. 3). We now outline the limitations of our model and discuss directions for future work.
First, the macroscopic Coulomb interaction $V_{\text{eff}}(\vec{q},\omega)$ is generally both wavevector and frequency dependent, and models of this function could be analyzed within a full Green’s function formalism for more quantitative analysis. Our model is best suited for the long-wavelength, static limit $V_{\text{eff}}(q\to 0,\omega\to 0)$, where we have additionally assumed that $V_{\text{eff}}$ is isotropic along the two in-plane directions of 2DES. In the context of engineered phase transitions, detailed knowledge of $V_{\text{eff}}(\vec{q},\omega)$ is required to predict which ordered ground state would be most stable. In the specific case of superconducting ground state, anisotropy of $V_{\text{eff}}$ in the plane of 2DES will affect whether the pairing is expected to be s-wave (as assumed here) or unconventional.
Second, the model for coupling between 2DES and PDT (Kittel model) in Eqs. (6),(7),(8) could be improved in two steps: (i) modeling of optic branch dispersion $\omega_{+}(q)$ (see Appendix C) and coupling $\Omega(q)$ to extend the present result for $\varepsilon^{f}_{z}(q)$ to finite $q$, and (ii) calculation of contribution from all types and branches of excitations in the PDT system [sidorkin2006domain], amounting to complete knowledge of $\varepsilon^{f}_{z}(\vec{q})$ in terms of $\omega_{i}^{L(T)}(\vec{q})$ and $\Omega_{i}^{L(T)}(\vec{q})$ for each branch $i$ and type - longitudinal (L) or transverse (T).
Third, the description of NC in $\text{PbTiO}_{3}$ based on the simple Kittel model may be inadequate when domain structures in experiments exhibit greater complexity (e.g., vortex textures [yadav2019]). Theoretical modeling of the collective excitations in such structures [yang2021condensation] is necessary to account for such details.
Fourth, the models used for 2DES and domain-provided NC are essentially zero-temperature theories. This could be extended by (i) using the Lindhard function at finite temperature [giuliani2008quantum] for $\Pi_{el}$ in Eq. (2) and (ii) incorporating temperature-dependent domain configurations, which would effectively make the NC response function temperature-dependent [pavlenko2021temperature].
Fifth, properties of the bare 2DES itself are expected to deviate significantly from the non-interacting description (used in this work) at low carrier densities. One well-known deviation is negative compressibility [eisenstein1992negative], in which case our assumption of a strictly positive $C_{q}$ from the non-interacting theory may not be a reliable approach. These corrections could be included in the present model by replacing $C_{q}$ with an expression that incorporates interaction effects [giuliani2008quantum].
Finally, we note that the superconducting transition temperature $T_{c}\sim\exp(-1/(\lambda-\mu^{*}))$ so that small errors in estimating the quantity $\lambda-\mu^{*}$ can cause large errors in estimating $T_{c}$. This makes it important to design experiments that can measure small values of $\lambda-\mu^{*}$ directly.
VI Acknowledgements
SD would like to thank Sayeef Salahuddin for alerting him to transient NC and for helpful discussions regarding negative permittivity in ferroelectrics. He would also like to thank Bhaskaran Muralidharan, Kerem Camsari and Zubin Jacob for their feedback on an earlier version of the manuscript. AS and PU gratefully acknowledge helpful discussions with Allan H. MacDonald on anisotropic gate screening.
This work was supported by the Purdue Research Foundation and the Blue Sky Research Program through the Purdue College of Engineering.
Appendix A Anisotropic gate screening effect with Negative Capacitance
We consider a two-dimensional electron system (2DES) encapsulated on one side by a dielectric medium with homogeneous positive static permittivity tensor, $\varepsilon^{d}=\mathrm{diag}(\varepsilon_{\perp}^{d},\varepsilon_{\perp}^{d},\varepsilon_{z}^{d})$, and on the other side by a negative capacitance (NC) medium described by $\varepsilon^{nc}=\mathrm{diag}(\varepsilon_{\perp}^{nc}>0,\varepsilon_{\perp}^{nc}>0,\varepsilon_{z}^{nc}<0)$, where $z$ denotes the stacking direction.
From Poisson’s equation, the electrostatic potential $\varphi$ in the two encapsulating regions takes the form:
$\begin{gathered}\varphi(q,z>0)=\alpha^{(d)}e^{\eta_{d}qz}+\beta^{(d)}e^{-\eta_{d}qz},\\ \varphi(q,z<0)=\alpha^{(nc)}e^{\eta_{nc}qz}+\beta^{(nc)}e^{-\eta_{nc}qz},\end{gathered}$ (14)
with anisotropy factors $\eta_{d}$ = $\sqrt{\varepsilon^{d}_{\perp}/\varepsilon^{d}_{z}}$ and $\eta_{nc}$ = $\sqrt{\varepsilon^{nc}_{\perp}/\varepsilon^{nc}_{z}}$. Metal gates are placed at distances $L_{d},L_{nc}$ above and below the $z=0$ plane where 2DES is located.
Imposing the boundary condition $\varphi(q,L_{d/nc})$ = 0 at the gates and enforcing continuity of $\varphi$ across the 2DES plane, we obtain:
$\begin{aligned} \varphi(q)\equiv\varphi(q,z=0)&=\alpha^{(nc)}(1-e^{-2\eta_{nc}qL_{nc}})\\ &=\beta^{(d)}(1-e^{-2\eta_{d}qL_{d}}).\end{aligned}$ (A2)
[…]
$\begin{aligned} &\alpha^{(nc)}\varepsilon_{z}^{nc}\partial_{z}(e^{\eta_{nc}qz}-e^{-2\eta_{nc}qL_{nc}}e^{-\eta_{nc}qz})\big|_{z=0}\\ &-\beta^{(d)}\varepsilon_{z}^{d}\partial_{z}(e^{-\eta_{d}qz}-e^{-2\eta_{d}qL_{d}}e^{\eta_{d}qz})\big|_{z=0}=\frac{e}{\varepsilon_{0}}.\end{aligned}$ (A3)
Solving for $\alpha^{(nc)}$ and substituting in Eq. (A2), the potential in 2DES is:
$\varphi(q)=\frac{e}{\varepsilon_{0}}\times\dfrac{1}{\varepsilon_{z}^{d}(q)\dfrac{\eta_{d}q}{\tanh(\eta_{d}qL_{d})}+\varepsilon_{z}^{nc}\dfrac{\eta_{nc}q}{\tanh(\eta_{nc}qL_{nc})}}.$ (15)
The corresponding environment contribution to the interaction energy, $V_{\text{env}}(q)=e\varphi(q)$, defines the effective background permittivity via
$V_{\text{env}}(q)=\frac{e^{2}}{2\varepsilon_{0}\varepsilon_{\text{env}}q}.$ (16)
Finally, we note that for anisotropic NC media $\eta_{nc}$ is purely imaginary, but the static quantities $\varphi(q)$, $V_{\text{env}}(q)$ (and $V_{\text{eff}}(q)$ in main text) remain strictly real.
Appendix B Stability of Negative Capacitance in MF2IM configuration
We consider a system where the dielectric is held at voltage $V_{d}$, the NC medium at $V_{nc}$, and the 2DES at $V_{q}$. For small charge fluctuations, the free energy of the system is
$\mathcal{F}=\frac{Q_{d}^{2}}{2C_{d}}+\frac{Q_{nc}^{2}}{2C_{nc}}+\frac{Q_{q}^{2}}{2C_{q}}-Q_{d}V_{d}-Q_{nc}V_{nc}-Q_{q}V_{q},$ (17)
subject to the constraint: $Q_{d}+Q_{nc}+Q_{q}=0$. Using this constraint to eliminate one variable, we obtain a reduced free energy $\tilde{\mathcal{F}}$, whose Hessian matrix is
[…]
The necessary and sufficient condition for $H$ to be positive definite (i.e., for both eigenvalues to be strictly positive) is
$\frac{1}{C_{d}C_{nc}}+\frac{1}{C_{nc}C_{q}}+\frac{1}{C_{q}C_{d}}>0.$ (18)
Assuming $C_{nc}<0$ and $C_{d},C_{q}>0$, this condition simplifies to the stability condition quoted in the main text as Eq. (5).
Appendix C Domain-provided Negative Capacitance model
In a ferroelectric with periodic domain texture (PDT), the ‘restoring force’ characterizing the response of domain walls to an external electric field can be calculated within the Kittel model of alternating ‘hard’ domains. The equilibrium domain width in this model [catalan2012domain] is:
$d=\sqrt{3.53\sqrt{\frac{\varepsilon^{f}_{\perp}}{\varepsilon^{f}_{z,hf}}}\xi\delta L_{f}},$ (19)
where $\xi=2\times(1+\varepsilon^{d}_{z}/\sqrt{\varepsilon^{f}_{\perp}\varepsilon^{f}_{z,hf}})$ is a parameter capturing electrostatic boundary conditions, $\delta\approx 1~\text{nm}$ is the domain wall thickness [luk2018prb], and $L_{f}$ is the ferroelectric thickness.
The PDT system has a ‘stiffness’ because there is an energy cost from long-range Coulomb forces when this system is displaced from equilibrium. The oscillation frequency corresponding to this stiffness is calculated using Fourier analysis:
$\begin{aligned} \omega_{\pm}^{2}&(q)=\frac{4P_{s}^{2}}{\pi\varepsilon_{0}\sqrt{\varepsilon^{f}_{\perp}\varepsilon^{f}_{z}}ML_{f}}\times\\ &\Bigg[\sum_{n=1}^{\infty}\textrm{ln}\bigg(1+\frac{\varepsilon^{f}_{\perp}}{\varepsilon^{f}_{z}}\frac{L_{f}^{2}}{d^{2}}\frac{1}{(2n-1)^{2}}\bigg)\big(1\pm\cos((2n-1)qd)\big)\\ &-\sum_{n=1}^{\infty}\textrm{ln}\bigg(1+\frac{\varepsilon^{f}_{\perp}}{\varepsilon^{f}_{z}}\frac{L_{f}^{2}}{d^{2}}\frac{1}{(2n)^{2}}\bigg)(1-\cos(2nqd))\Bigg],\end{aligned}$ (C2)
where the two branches ($\pm$) of excitation for this system are analogous to the acoustic and optic branches in crystals with a two-atom basis. In this work we restrict attention to the ++ (‘optic’) branch in the long-wavelength limit, $q\to 0$, so $\omega_{0}$ in the main text is just $\omega_{+}(q=0)$. Within this model, the calculated negative permittivity $\varepsilon^{f}_{z}$ has a value of about 60-60 for $L_{f}=4~\text{nm}$.

### v2

@@
 Abstract
-It is known that the many-body ground state of a two-dimensional electron system can be tuned through Coulomb engineering by controlling the permittivity of the surrounding media.
-However, permittivities are traditionally restricted to positive values.
-In this paper we argue that the negative capacitance effect demonstrated in appropriately engineered structures can open new vistas in Coulomb engineering.
-Negative permittivities transform the natural repulsive interaction of electrons into an attractive one raising the intriguing possibility of a superconducting ground state.
+The many-body ground state of a two-dimensional electron system can be tuned by Coulomb engineering through control of the dielectric environment.
+However, in conventional dielectrics the static permittivity is restricted to positive values, limiting the accessible interaction regimes.
+Here we argue that the negative capacitance demonstrated in appropriately engineered structures can open new vistas for Coulomb engineering.
+The associated negative permittivity could transform the natural repulsive interaction of electrons into an attractive one, raising the intriguing possibility of nontrivial ground states, including superconductivity.
 Using models of two-dimensional electron systems with linear and parabolic dispersion relations coupled to environments with negative capacitance, we estimate the strength and sign of the engineered Coulomb interaction and outline parameter regimes that could stabilize correlated electronic phases.
 preprint: APS/123-QED
-I Introduction
-Coulomb interactions between electrons govern a variety of many-body phenomena, from stabilizing interacting phases to shaping the electrical and optical responses of matter.
+Introduction.—Coulomb interactions between electrons govern a variety of many-body phenomena, from stabilizing interacting phases to shaping the electrical and optical responses of matter.
 Consequently, several approaches have been developed to highlight the effects driven by Coulomb interactions, bringing them to the forefront of condensed matter physics.
-For example, one prominent approach [bistritzer2011moire] focuses on revealing Coulomb interactions by suppressing competing kinetic energies in materials engineered to host flat energy bands.
-This flat-band engineering technique has successfully stabilized interacting phases ranging from superconductivity [cao2018, lu2019superconductors, balents2020superconductivity, oh2021] to the fractional quantum anomalous Hall state [sharpe2019qahe, serlin2020qahe, li2021qahe] in two-dimensional materials [kennes2021moirereview].
-However, because this approach does not directly modify Coulomb interactions, the resulting interacting phases are typically stabilized only at low temperatures.
-A complementary strategy—known as Coulomb engineering—enables the direct tuning of Coulomb interactions by designing heterostructures with tailored electromagnetic environments [raja2017, song2018].
-Although Coulomb engineering has successfully altered the electrical and optical responses of materials [jang2008graph, hwang2012fvr, yu2013qcap, kajino2019coulomb, whelan2020fvr], it remains challenging to tune Coulomb interactions over a wide range and to design interaction-driven phases of matter within this framework.
-This limitation arises because conventional electromagnetic environments explored so far are largely constrained by the requirement that their static dielectric function (permittivity) must remain positive and be greater than that of vacuum.
-This work is motivated by recent experimental demonstrations of Negative Capacitance (NC) [zubko2016, yadav2019, das2021local].
-We propose to modify the structures in those works to embed two-dimensional electron systems (2DES) such that the normally repulsive interaction between electrons is rendered attractive, thereby unveiling a new regime of unconventional Coulomb engineering.
-The space of applications unlocked by this extended tuning of Coulomb interactions spans both quantitative modifications and qualitative transformations.
-On the quantitative side, one can expect changes to transport properties and optical response in the normal state of 2DES, potentially exceeding limits of conventional Coulomb engineering [jang2008graph, hwang2012fvr, yu2013qcap, raja2017, song2018, kajino2019coulomb, whelan2020fvr].
-On the qualitative side, entirely new phases of matter may emerge [van2023cemott], especially once the customarily repulsive electron–electron interactions are rendered attractive.
-Of particular interest is superconductivity, which requires the formation of paired electronic states [bcs1957].
-We provide estimates of pairing strength through the dimensionless parameter $\lambda-\mu^{*}$, adapted from electron–phonon theory [tolmachev1961, morelanderson, colemanbook], and find that values $\gtrsim 0.1$ are achievable in realistic regimes.
-Crucially, this parameter is highly tunable within our proposal, suggesting that NC-enabled Coulomb engineering could provide a new route toward stabilizing superconductivity and other ordered electronic phases, potentially even at elevated temperatures.
-II Central Idea
-To illustrate the main concept, we begin with the idealized structure shown schematically in Fig.
+For example, one prominent approach [5] focuses on revealing Coulomb interactions by suppressing competing kinetic energies in materials engineered to host flat energy bands.
+This technique has successfully stabilized interacting phases ranging from superconductivity [6, 37, 3, 42] to the fractional quantum anomalous Hall state [54, 53, 35] in two-dimensional materials [28].
+A complementary strategy—known as Coulomb engineering—enables the direct tuning of Coulomb interactions by designing heterostructures with tailored electromagnetic environments [49, 56].
+Although Coulomb engineering has successfully altered the electrical and optical responses of materials [25, 22, 63, 27, 60], it remains challenging to tune Coulomb interactions over a wide range and to design interaction-driven phases of matter within this framework.
+This limitation arises because previously explored electromagnetic environments are largely restricted to positive static permittivities.
+However, on fundamental grounds, environments with local static negative permittivity are permissible [11], raising the question of whether present-day Coulomb-engineered structures can be extended beyond this constraint.
+Here we answer this question in the affirmative by proposing and studying novel Coulomb-engineered structures motivated by recent demonstrations of ferroelectric negative capacitance (NC) [65, 61, 10].
+Specifically, we propose to modify the structures in those works to embed two-dimensional electron systems (2DES) such that the normally repulsive interaction between electrons is rendered attractive, thereby unveiling a new regime of unconventional Coulomb engineering.
+On the one hand, this extended tuning of Coulomb interactions can change the transport properties and optical response in the normal state of 2DES, potentially exceeding limits of conventional Coulomb engineering [25, 22, 63, 49, 56, 27, 60].
+On the other hand, entirely new phases of matter may emerge [58], especially once the customarily repulsive electron–electron interactions are rendered attractive.
+While predicting the ground state is beyond the scope of the present work, a state of particular interest with attractive electron interactions is superconductivity, which requires the formation of paired electronic states [4].
+Motivated by this, we estimate the pairing strength in a model platform where attractive interactions are mediated by negative capacitance arising from the motion of ferroelectric domain walls [39].
+Within this model, the induced attraction is retarded by the ferroelectric domain wall response, whose characteristic frequencies typically lie well below electronic energy scales, leading to a separation of timescales analogous to the phonon-mediated pairing.
+We find that the resulting dimensionless pairing parameter $\lambda-\mu^{*}$, adapted from electron–phonon theory [57, 41, 8], can reach values $\gtrsim 0.1$ in realistic experimental regimes.
+Crucially, by balancing positive and negative capacitances, this parameter is highly tunable within our proposal, suggesting that NC-enabled Coulomb engineering could provide a new route towards designing superconductivity and other ordered electronic phases.
+More broadly, there has recently been growing interest in cavity-based materials engineering [51, 36], where the collective response of specially designed cavities is leveraged to modify the ground state of solid-state materials.
+Our results indicate that the collective response of ferroelectrics stabilized in the negative-capacitance regime could offer a useful resource for this paradigm.
+Central Idea.—To illustrate the main concept, we begin with the idealized structure shown schematically in Fig.
 1(a) (connections to real material platforms are made in later sections).
@@
 We model the DE and NC regions with diagonal permittivity tensors $\varepsilon^{nc(d)}=\text{diag}(\varepsilon^{nc(d)}_{\perp},\varepsilon^{nc(d)}_{\perp},\varepsilon^{nc(d)}_{z})$, where $z$ denotes the stacking direction and the superscripts $nc$ and $d$ represent NC and DE regions, respectively.
-In the 2DES plane, the static Coulomb interaction $V_{\text{eff}}$ as a function of in-plane wavenumber $q$ is given by
-$V_{\text{eff}}(q)=\frac{e^{2}}{2\varepsilon_{0}q}\cdot\frac{1}{\varepsilon_{\text{eff}}(q)}=\frac{v_{q}}{\varepsilon_{\text{eff}}(q)},$ (1)
-where the effective dielectric function $\varepsilon_{\text{eff}}(q)$ is obtained self-consistently as
+In the 2DES plane, the static Coulomb interaction $V_{\text{eff}}$ as a function of in-plane wavenumber $q$ is given by:
+$V_{\mathrm{eff}}(q)=\frac{v_{q}}{\varepsilon_{\mathrm{eff}}(q)},\qquad v_{q}\equiv\frac{e^{2}}{2\varepsilon_{0}q}.$ (1)
+We model the effective dielectric function $\varepsilon_{\mathrm{eff}}(q)$ using linear response polarization functions as: $\varepsilon_{\mathrm{eff}}(q)=1-v_{q}(\Pi_{nc}(q)+\Pi_{d}(q)+\Pi_{el}(q))$.
+The environment dielectric function $\varepsilon_{\mathrm{env}}(q)\equiv 1-v_{q}(\Pi_{nc}(q)+\Pi_{d}(q))$ can be derived from Poisson’s equation (see Appendix A) and reflects the ‘gate screening’ effect:
 […]
-  • The expression for $\varepsilon_{\text{env}}(q)$ reflects the “anisotropic gate screening” effect, derived from Poisson’s equation (see Appendix A).
-  • For the electron polarization function $\Pi_{el}$, we use the static limit of the Lindhard function for a 2DES [giuliani2008quantum], which reduces to the Thomas-Fermi (TF) form for wavenumber $q<2k_{F}$—the range relevant for this work.
 Figure 1: Central idea.
-(a) System schematic - a two-dimensional electron system (2DES) surrounded by dielectric (conventional) and negative capacitance (NC) media, with respective thicknesses $L_{d}$ and $L_{nc}$.
-When the NC material is a ferroelectric, this configuration will be referred to as a Metal-Ferroelectric-2DES-Insulator-Metal (MF2IM) structure.
-(b) Tunability of the engineered 2DES Coulomb interaction $V_{\text{eff}}$ in the long-wavelength approximation as per Eq.
+(a) System schematic: a two-dimensional electron system (2DES) surrounded by dielectric (conventional) and negative-capacitance (NC) media, with respective thicknesses $L_{d}$ and $L_{nc}$.
+When the NC material is a ferroelectric, this structure is referred to as Metal-Ferroelectric-2DES-Insulator-Metal (MF2IM) configuration.
+(b) Long-wavelength part of engineered 2DES Coulomb interaction $V_{\text{eff}}$ as per Eq.
 (3), shown as a function of geometric capacitances $C_{nc}$ and $C_{d}$, and quantum capacitance $C_{q}$.
-The dashed black line indicates the limit of conventional Coulomb engineering ($C_{nc}>0$, region I).
-Notably, $V_{\text{eff}}$ can become negative (region III) if $C_{nc}$ is sufficiently negative.
-Long-wavelength limit: Introducing dimensionless anisotropy factors $\eta_{nc}=\sqrt{\varepsilon^{nc}_{\perp}/\varepsilon^{nc}_{z}}$ and $\eta_{d}=\sqrt{\varepsilon^{d}_{\perp}/\varepsilon^{d}_{z}}$, we illustrate the central idea of NC-enabled unconventional Coulomb engineering by first considering the limit where $q\eta_{nc}L_{nc}\ll 1$ and $q\eta_{d}L_{d}\ll 1$.
-In this regime, the wavelengths associated with low-energy scattering of electrons in the 2DES are large compared to the gate distances.
-The assumption of $q$-independent permittivities in this section additionally requires that this wavelength is large compared to characteristic length scales of the polarizable DE and NC media (elaborated in later sections).
-Under these conditions, Eq.
-(2) reduces to a simple form:
+The black dashed line indicates the limit of conventional Coulomb engineering ($C_{nc}>0$, region I).
+Region II is new but expected to be unstable in the MF2IM configuration.
+Notably, $V_{\text{eff}}$ can become negative in region III, where thermodynamic stability is satisfied as per Eq.
+(4).
+Introducing dimensionless anisotropy factors $\eta_{nc}=\sqrt{\varepsilon^{nc}_{\perp}/\varepsilon^{nc}_{z}}$ and $\eta_{d}=\sqrt{\varepsilon^{d}_{\perp}/\varepsilon^{d}_{z}}$, we illustrate the central idea of NC-enabled unconventional Coulomb engineering by first considering the limit where $q\eta_{nc}L_{nc}\ll 1$ and $q\eta_{d}L_{d}\ll 1$.
+In this regime, the wavelengths associated with low-energy scattering of electrons in the 2DES are large compared to the gate distances and characteristic length scales of the polarizable DE and NC media (elaborated in later sections).
+For such large wavelengths ($q\to 0$), Eqs.
+(1)–(2) reduce to a simple form:
 $V_{\text{eff}}\cong\frac{e^{2}}{C_{nc}+C_{d}+C_{q}},$ (3)
-where $C_{nc}$ and $C_{d}$ are geometric capacitances per unit area, and $C_{q}$ is the quantum capacitance of the 2DES:
-$C_{nc}=\frac{\varepsilon_{0}\varepsilon^{nc}_{z}}{L_{nc}}\hskip 5.69046pt;\hskip 5.69046ptC_{d}=\frac{\varepsilon_{0}\varepsilon^{d}_{z}}{L_{d}}\hskip 5.69046pt;\hskip 5.69046ptC_{q}=2\varepsilon_{0}k_{\text{TF}}.$ (4)
-Throughout this work, we set the ‘background’ dielectric constant of 2DES to unity for simplicity.
-Eq.
-(3) encapsulates our central thesis: the inclusion of $C_{nc}<0$ enables the design of structures where $V_{\text{eff}}$ can be negative and engineered to have large magnitude values through careful matching of the constituent capacitances.
+where $C_{nc}\equiv\varepsilon_{0}\varepsilon^{nc}_{z}/{L_{nc}}$, $C_{d}\equiv\varepsilon_{0}\varepsilon^{d}_{z}/{L_{d}}$ are geometric capacitances per unit area, and $C_{q}\equiv-e^{2}\Pi_{el}(q\to 0)$ is the quantum capacitance of the 2DES.
+The expression for $C_{q}$ reflects the compressibility sum rule [16].
+For conventional dielectrics, $C_{d}>0$, whereas $C_{nc}$ can be negative for materials operating in the NC regime.
+Throughout this work we focus on the case of $C_{q}>0$, which may be violated in certain low-density regimes of a 2DES due to exchange and correlation effects [13].
 Before discussing the scope of this result, we first present a stability condition which constrains the values $C_{nc}$ can assume in the proposed configuration.
-Structural Stability: The NC state of a material in isolation is thermodynamically unstable, but as argued in Ref.
-[salahuddindatta2008], it can be locally stabilized in a constituent layer provided the composite structure is stable against charge fluctuations.
+The NC state of a material in isolation is thermodynamically unstable, but as argued in Ref.
+[52], it can be locally stabilized in a constituent layer provided the composite structure is stable against charge fluctuations.
 We refer to the structure in Fig.
-1(a) as Metal-Ferroelectric-2DES-Insulator-Metal (MF2IM) configuration (specifying NC material as ferroelectric to connect with notation used in literature [park2019modeling, luk2022ferroelectric]).
-The necessary and sufficient condition for stability of NC in the MF2IM configuration is (assuming $C_{nc}<0$ and $C_{d},C_{q}>0$, see Appendix B for details):
+1(a) as Metal-Ferroelectric-2DES-Insulator-Metal (MF2IM) configuration (specifying NC material as ferroelectric to connect with notation used in literature [44, 38]).
+The necessary and sufficient condition for stability of NC in the MF2IM configuration is (assuming $C_{nc}<0$, and $C_{d},C_{q}>0$, see Appendix B for details):
 $C_{nc}+C_{d}+C_{q}<0.$ (4)
-Setting $C_{q}=0$ recovers the familiar stability condition for the widely studied Metal-Ferroelectric-Insulator-Metal (MFIM) structure [salahuddindatta2008, park2019modeling, iniguez2019ferroelectric].
+Setting $C_{q}=0$ recovers the familiar stability condition for the widely studied Metal-Ferroelectric-Insulator-Metal (MFIM) structure [52, 44, 24].
 In the opposite limit of $C_{q}\to\infty$, the MF2IM structure decouples into two separate capacitors (assuming the 2DES is connected to a charge reservoir), and NC cannot be stabilized in this case.
-Scope: Fig.
-1(b) illustrates the scope of Eq.
-(3) by treating the capacitances as free variables.
-In conventional Coulomb engineering where all geometric capacitances are positive, the tunability of $V_{\text{eff}}$ is limited to region I (assuming $C_{q}>0$).
-The inclusion of $C_{nc}<0$ opens up new regimes, indicated as regions II and III.
-Most strikingly, in region III the sign of $V_{\text{\text{eff}}}$ can be reversed while still satisfying the stability condition in Eq.
-(5), transforming the naturally repulsive electron-electron interaction into an effective attraction.
+Fig.
+1(b) illustrates the first main result of this work by treating the capacitances as free variables.
+In conventional Coulomb engineering, where all capacitances are positive, the tunability of $V_{\text{eff}}$ is limited to region I.
+The inclusion of $C_{nc}<0$ opens additional regimes, labeled regions II and III.
+According to Eq.
+(4), region II is unstable against charge fluctuations.
+If the system is tuned into this region—for example, by varying $C_{q}$—the charge distribution will reconfigure to restore stability.
+This behavior may provide a new diagnostic tool for probing NC materials.
+In contrast, region III remains stable while reversing the sign of $V_{\mathrm{eff}}$, converting the intrinsically repulsive electron–electron interaction into an effective attraction.
+Moreover, as per Eq.
+(3), the magnitude of this attractive $V_{\mathrm{eff}}$ can be engineered to reach large values through careful matching of the constituent capacitances.
+In the remainder of this work, we focus on region III, emphasizing model experimental platforms capable of realizing this regime.
+Model Platform.— Here, we focus on ferroelectrics as the specific choice of NC material.
+Within the growing literature on negative capacitance, one distinguishes transient NC [30, 29, 19, 17, 31, 47], which occurs during polarization switching and reflects nonequilibrium order-parameter relaxation, from stabilized NC [2, 15, 65, 20, 61, 10, 47, 48], corresponding to a quasi-static regime with negative permittivity.
+NC systems are further classified by material class—perovskites [61], oxides [18], van der Waals materials [59]—and by ferroic order, including ferroelectrics  [24] and antiferroelectrics [47, 21].
+Here we focus on stabilized NC, distinguishing extrinsic NC [65, 61, 10], associated with multidomain configurations and domain-wall motion [26], from intrinsic NC [18], which appears independent of domain structure.
+While the general framework presented in Fig.
+1 is agnostic to the underlying mechanism, our material-specific calculations (Figs.
+2 and 3) employ a domain-based model of extrinsic NC, which we now detail for stabilized NC in perovskite ferroelectrics, exemplified by $\text{PbTiO}_{3}$.
+This choice is motivated by substantial evidence from experiments [65, 61, 10] and corresponding theoretical explanation [39, 40] providing a framework for modeling near-equilibrium NC.
+In ferroelectric $\text{PbTiO}_{3}$ ($\varepsilon^{nc}_{z(\perp)}\to\varepsilon^{f}_{z(\perp)}$, $L_{nc}\to L_{f}$), the polarization response is highly anisotropic, with $\varepsilon^{f}_{z}<0$ and $\varepsilon^{f}_{\perp}>0$.
+Starting from a periodic domain texture (PDT) as the equilibrium polarization configuration, the static negative permittivity $\varepsilon^{f}_{z}$ is explained as an overscreening effect due to the role of depolarization field in this system [39, 40].
+Assuming small deviations from equilibrium, the PDT dynamics can be described by an oscillator model for domain wall displacements, where $P$ is the electric polarization and $E_{\text{tot}}$ is the total electric field (external + depolarization):
+$\ddot{P}(t)+(\omega_{0}^{2}-\Omega^{2})P(t)=\Omega^{2}\varepsilon_{0}\varepsilon^{f}_{z,hf}E_{\text{tot}}(t).$ (5)
+The restoring force originates from Coulomb energy associated with excess surface charges when the PDT is displaced from equilibrium.
+Here, $\omega_{0}$ is the characteristic oscillation frequency of the PDT system and $\Omega$ represents the strength of coupling to external electric field.
+We set the damping [39] to zero, as we focus on the static response in this work.
+In a more detailed model, the frequency parameters should exhibit dispersion $\omega_{0}(q),\Omega(q)$ with additional indices for excitation type (longitudinal vs transverse) and branch (acoustic vs optic).
+The expressions which we use to calculate $\omega_{0},\Omega$ in this work (see Appendix C) account for only the longitudinal optic branch at zero wavenumber, based on the expectation that this mode contributes dominantly to the NC effect in the long-wavelength limit (defined below).
+We refer to this as the ‘single-mode’ approximation.
 Figure 2: Critical length scales.
-(a) Definition of two regimes based on three length scales.
-The wavelength of low-energy electron scattering in the 2DES relative to $2d$ (twice the domain width) in ferroelectric periodic domain texture (PDT) governs whether dispersion $\varepsilon^{f}_{z}(q)$ is important in extrinsic NC systems ($2d$ should be replaced by an appropriate lattice-related length scale in the case of intrinsic NC) and relative to $|\eta_{f}|L_{f}$ determines whether the linear approximation of gate screening effect is valid.
+(a) The wavelength of low-energy electron scattering in the 2DES, $(2k_{F})^{-1}$, relative to $2d$ (twice the domain width) in ferroelectric periodic domain texture (PDT) governs whether dispersion $\varepsilon^{f}_{z}(q)$ is important in extrinsic NC case, and relative to $|\eta_{f}|L_{f}$ (scaled ferroelectric thickness) determines whether the linear approximation of gate screening effect is valid.
+In the intrinsic NC case, $2d$ should be replaced by an appropriate lattice-related length scale.
 (b), (c) Illustration of effective Coulomb interaction energy $V_{\text{eff}}$ as a function of wavenumber $q$ as per Eqs.
-(1) and (2), shown for linear and parabolic 2DES with typical $v_{F}$ and $m^{*}$ values.
-In both cases, 2DES carrier density $n$ was varied to access the long-wavelength (low $n=10^{15}/\text{m}^{2}$) and short-wavelength (high $n=10^{17}/\text{m}^{2}$) regimes.
-Note that changing $L_{f}$ (fixed at $4~\text{nm}$ for this figure) is an alternate way to access these regimes.
-III Model details
-In this section, we provide details of the models we use to estimate parameters appearing in Eq. (2), treating the 2DES carrier density $n$ and the NC material thickness $L_{nc}$ as tunable variables.
-NC material: Within the growing literature on negative capacitance, it is useful to distinguish between transient and stabilized NC.
-Transient NC, such as that reported in [khan2015transient, khan2017transient, hoffmann2018transient, hoffmann2019unveiling, kim2019transient, qiao2021both], occurs during polarization switching and is therefore a non-equilibrium associated with dynamical relaxation of the order parameter.
-In contrast, stabilized NC [appleby2014stable, gao2014stable, zubko2016, hoffmann2018stabilization, yadav2019, das2021local, qiao2021both, qiao2024stable] refers to the quasi-static, near-equilibrium regime, where the system exhibits static negative permittivity.
-Beyond this dynamical classification, NC systems can also be categorized by material class— perovskite (such as [yadav2019]), oxide (such as [hoffmann2022intrinsic]), van der Waals (such as [wang2019vdw])—and by the underlying ferroic order, with both ferroelectrics [iniguez2019ferroelectric] and antiferroelectrics [qiao2021both, hoffmann2022antiferroelectric] being reported to exhibit NC behavior.
-In the remainder of this work we restrict attention to stabilized NC, although the idea outlined in the previous section could be explored in the transient case as well.
-Within stabilized NC, two classes of systems are often distinguished: (i) extrinsic NC [zubko2016, yadav2019, das2021local], where the phenomenon is associated with multi-domain configurations and domain wall motion [junquera2023topological], and (ii) intrinsic NC [hoffmann2022intrinsic], where the effect appears to be independent of specific domain configurations.
-The theory developed in the previous section and summarized in Fig.
-1 is formulated without assuming either classification, but we adopt a domain-provided model for extrinsic NC when performing calculations for specific material systems (results in Figs.
-2 and 3).
-We now provide details about the domain-provided model for stabilized NC in perovskite ferroelectric $\text{PbTiO}_{3}$.
-This choice is motivated by substantial evidence from experiments [zubko2016, yadav2019, das2021local] and corresponding theoretical explanation [luk2018prb, luk2019harnessing] providing a framework for modeling near-equilibrium NC.
-In ferroelectric $\text{PbTiO}_{3}$ ($\varepsilon^{nc}_{z(\perp)}\to\varepsilon^{f}_{z(\perp)}$, $L_{nc}\to L_{f}$), the polarization response is highly anisotropic, with $\varepsilon^{f}_{z}<0$ and $\varepsilon^{f}_{\perp}>0$.
-Starting from a periodic domain texture (PDT) as the equilibrium polarization configuration, the static negative permittivity $\varepsilon^{f}_{z}$ is explained as an overscreening effect due to the role of depolarization field in this system [luk2018prb, luk2019harnessing].
-Assuming small deviations from equilibrium, the PDT dynamics can be described by an oscillator model for domain wall displacements, where $P$ is the electric polarization and $E_{\text{tot}}$ is the total electric field (external + depolarization):
-$\ddot{P}(t)+\Gamma\dot{P}(t)+(\omega_{0}^{2}-\Omega^{2})P(t)=\Omega^{2}\cdot\varepsilon_{0}\varepsilon^{f}_{z,hf}E_{\text{tot}}(t).$ (6)
-Here, $\Gamma$ is the damping constant for domain wall motion.
-The restoring force originates from Coulomb energy associated with excess surface charges when the PDT is displaced from equilibrium.
-Fourier analysis of the domain structure within Kittel model [kittel1946, junquera2023topological] yields the following expressions for frequency parameters [sidorkin2006domain, luk2018prb]:
-$\begin{aligned} \omega_{0}^{2}=\frac{8P_{s}^{2}}{\pi\varepsilon_{0}\sqrt{\varepsilon^{f}_{\perp}\varepsilon^{f}_{z,\text{hf}}}ML_{f}}&\mathrm{ln}\Bigg(\mathrm{cosh}\Bigg(\sqrt{\frac{\varepsilon^{f}_{\perp}}{\varepsilon^{f}_{z,\text{hf}}}}\frac{\pi}{2}\frac{L_{f}}{d}\Bigg)\Bigg)\\ \Omega^{2}=&\dfrac{4P_{s}^{2}}{\varepsilon_{0}\varepsilon^{f}_{z,\text{hf}}Md}.\end{aligned}$ (7)
-Here, $\omega_{0}$ is the characteristic oscillation frequency of the PDT system and $\Omega$ represents the strength of coupling to external electric field.
-In a more detailed model, these frequency parameters should exhibit dispersion $\omega_{0}(q),\Omega(q)$ with additional indices for excitation type (longitudinal vs transverse) and branch (acoustic vs optic).
-The expressions in Eq.
-(7) account only for the longitudinal optic branch at zero wavenumber (see Appendix C), based on the expectation that this mode contributes dominantly to the NC effect in the long-wavelength limit.
-We refer to this as the ‘single-mode’ approximation.
-Parameters for ferroelectric $\text{PbTiO}_{3}$ are chosen as follows [luk2018prb] : $P_{s}=0.65~\text{Cm}^{-2}$ is the spontaneous polarization, and $M$ is the domain wall mass per unit area estimated using interpolation formula $M=1.3\sqrt{L_{f}[\text{nm}]}\times 10^{-9}~\text{kgm}^{-2}$ [luk2018prb, zhangponomareva].
-The equilibrium domain width $d$ is calculated using the Kittel model (explicit formula in Appendix C).
-$\varepsilon^{f}_{z,hf}=100$ is the “high-frequency” dielectric constant which captures the background polarizability of $\text{PbTiO}_{3}$ arising from other polarization mechanisms in the material, and $\varepsilon^{f}_{\perp}=30$.
+(1)–(2), for linear and parabolic 2DES with typical parameter values $v_{F}=1\times 10^{6}~\text{m}/\text{s}$ and $m^{*}=0.1m_{e}$ respectively.
+Here, 2DES carrier density $n$ was varied to access the long-wavelength (low $n=10^{15}/\text{m}^{2}$) and short-wavelength (high $n=10^{17}/\text{m}^{2}$) regimes for fixed $L_{f}=4~\text{nm}$.
 The static linear response dielectric function corresponding to Eq. (5) is:
-$\varepsilon^{f}_{z}\equiv\varepsilon^{f}_{z}(q=0,\omega=0)=\ \varepsilon^{f}_{z,\text{hf}}\cdot\frac{\omega_{0}^{2}}{\omega_{0}^{2}-\Omega^{2}}.$ (8)
-Since $\omega_{0}<\Omega$ for ferroelectric $\text{PbTiO}_{3}$ in this model, the above expression yields a negative $\varepsilon^{f}_{z}$.
-Dielectric: For the conventional dielectric part of environment (see Fig.
+$\varepsilon^{f}_{z}\equiv\varepsilon^{f}_{z}(q=0,\omega=0)=\frac{\varepsilon^{f}_{z,\mathrm{hf}}\,\omega_{0}^{2}}{\omega_{0}^{2}-\Omega^{2}}.$ (6)
+Here $\varepsilon^{f}_{z,hf}>0$ denotes the high-frequency dielectric constant (see Appendix C for more details).
+Since $\omega_{0}<\Omega$ for ferroelectric $\text{PbTiO}_{3}$ in this model, the expression in Eq.
+(6) yields a negative $\varepsilon^{f}_{z}$.
+For the conventional dielectric part of environment (see Fig.
 1(a)), we adopt values for hexagonal boron nitride (hBN) as a representative ‘low-k’ dielectric.
-Specifically, we set $\varepsilon^{d}_{z}=3.4$, $\varepsilon^{d}_{\perp}=6.86$ based on values reported in literature [ohbahbn, laturiahbn, Pierrethbn, kimhbn].
+Specifically, we set $\varepsilon^{d}_{z}=3.4$, $\varepsilon^{d}_{\perp}=6.86$ based on values reported in literature [43, 34, 46, 32].
 In Fig.
@@
 The dielectric thickness is fixed at $L_{d}=4~\text{nm}$ in all cases.
-2DES: The bare 2DES (in vacuum, $\varepsilon_{\text{env}}=1$) is modeled as a single-valley non-interacting electron gas, with either linear or parabolic low-energy dispersion.
+The bare 2DES (in vacuum, $\varepsilon_{\text{env}}=1$) is modeled as a single-valley non-interacting electron gas, with either linear or parabolic low-energy dispersion.
 The two cases are characterized respectively by a Fermi velocity $v_{F}$, or an effective mass $m^{*}$.
 Spin and valley degeneracies are set to $g_{s}=2$, $g_{v}=1$ respectively.
-The quantum capacitance $C_{q}$ and Thomas-Fermi screening wavenumber $k_{\text{TF}}$ are
-$\begin{aligned} \text{Linear:}\hskip 5.69046ptC_{q}=2\varepsilon_{0}k_{\text{TF}}&=\frac{e^{2}k_{F}}{\pi\hbar v_{F}},\\ \text{Parabolic:}\hskip 5.69046ptC_{q}=2\varepsilon_{0}k_{\text{TF}}&=\frac{e^{2}m^{*}}{\pi\hbar^{2}}.\end{aligned}$ (9)
-Note that in the linear case $C_{q}$ and $k_{\text{TF}}$ scale with carrier density as $E_{F}$ ($\propto\sqrt{n}$). In the parabolic case, $C_{q}$ and $k_{TF}$ are independent of $n$.
-IV Results
-In this section, we apply the models described above to identify critical length and energy scales in the composite structure and estimate the strength of engineered Coulomb interactions in the low-energy regime.
-Length scales: Let us first note that the quantity $\eta_{nc}\equiv\sqrt{\varepsilon^{nc}_{\perp}/\varepsilon^{nc}_{z}}$ is imaginary for an anisotropic NC material like $\text{PbTiO}_{3}$ which has $\varepsilon^{f}_{z}<0~,~\varepsilon^{f}_{\perp}>0$.
-Consequently the tanhtanh functions in Eq.
-(2) become tantan functions ($\tanh(ix)=i\tan(x)$):
-$\varepsilon_{\text{env}}(q)=\frac{1}{2}\bigg[\frac{|\eta_{f}|\varepsilon^{f}_{z}}{\tan\big(q\hskip 0.85355pt|\eta_{f}|L_{f}\big.)}+\frac{\eta_{d}\varepsilon^{d}_{z}}{\tanh\big(q\hskip 1.42271pt\eta_{d}L_{d}\big.)}\bigg],$ (10)
+For the electronic screening, we use the Thomas-Fermi (TF) model and set $v_{q}\Pi_{el}=-k_{\mathrm{TF}}/q$, where the TF screening wavenumber $k_{\text{TF}}$ [9] is given by:
+$k_{\mathrm{TF}}=\frac{e^{2}}{\pi\hbar}\times\begin{cases}k_{F}/v_{F},&\text{Linear},\\ m^{*}/\hbar,&\text{Parabolic}.\end{cases}$ (7)
+Neglecting exchange and correlation effects, the compressibility sum rule is satisfied with $C_{q}=2\varepsilon_{0}k_{\mathrm{TF}}$.
+We set the ‘background’ dielectric constant of 2DES to unity for simplicity.
+Results.—We begin by analyzing the Fourier-transformed Coulomb interaction in the model platform introduced above.
+To this end, we first note that the quantity $\eta_{nc}\equiv\sqrt{\varepsilon^{nc}_{\perp}/\varepsilon^{nc}_{z}}$ is imaginary for an anisotropic NC material like $\text{PbTiO}_{3}$ which has $\varepsilon^{f}_{z}<0~,~\varepsilon^{f}_{\perp}>0$.
+Consequently the tanhtanh function in Eq.
+(2) becomes a tantan function ($\tanh(ix)=i\tan(x)$):
+$\varepsilon_{\mathrm{env}}(q)=\frac{1}{2}\bigg[\frac{|\eta_{f}|\varepsilon^{f}_{z}}{\tan\big(q\hskip 0.85355pt|\eta_{f}|L_{f}\big)}+\frac{\eta_{d}\varepsilon^{d}_{z}}{\tanh\big(q\hskip 1.42271pt\eta_{d}L_{d}\big)}\bigg],$ (8)
 where $|\eta_{f}|$ denotes the complex magnitude, and we have assumed that the dielectric part is conventional ($\varepsilon^{d}_{z},\varepsilon^{d}_{\perp}>0$).
@@
 For a system with isotropic NC ($\varepsilon^{f}_{z}<0$, $\varepsilon^{f}_{\perp}<0$) $\eta_{f}$ is real-valued, the corresponding gate screening function remains tanhtanh, and $V_{\text{eff}}(q)$ is not periodic.
-In addition to $|\eta_{f}|L_{f}$, two important length scales in the composite structure are $2k_{F}$ (low-energy scattering shell in the 2DES) and a characteristic length scale corresponding to the NC phenomenon, taken to be the domain width $d$ of PDT in the extrinsic case.
-We define two distinct regimes based on the relative values of these length scales: long-wavelength and short-wavelength.
-In the long-wavelength limit, two simplifications emerge: wavelengths of electron scattering in 2DES are too large to resolve microscopic details on the scale of domain width $d$ (so that a simplified model with constant $\varepsilon^{nc}(q)$ suffices), and gate screening reduces to its linear limit since the gates see an effectively uniform charge distribution in the 2DES.
-Expressed as mathematical conditions:
-  • In the long-wavelength regime, $2k_{F}d\ll 1$ and $2k_{F}|\eta_{f}|L_{f}\ll 1$, so that the single-mode approximation discussed in previous section is a reasonable starting point, and the simplified description of Eq.
-(3) is applicable [see Fig.
-2(b)].
-  • In the short-wavelength regime, $2k_{F}d\gg 1$ and $2k_{F}|\eta_{f}|L_{f}\gg 1$ , indicating that it is important to model dispersion $\varepsilon^{f}_{z}(q)$ (single-mode approximation may not necessarily be reliable) and gate screening modifies the $q$-space structure of $V_{\text{eff}}$ significantly [see Fig.
-2(c)].
-Figs.
-2(b),(c) demonstrate how the two length scale regimes are accessed using 2DES carrier density $n$ as the control parameter in a model system with ferroelectric $\text{PbTiO}_{3}$ as the NC material.
-$V_{\text{eff}}$ as a function of wavenumber $q$ is shown.
-Ferroelectric thickness $L_{f}$ could be varied at fixed $n$ to achieve the same effect.
-The long-wavelength limit is accessed in the case of low carrier density and small thickness.
-Energy scales: The low-energy picture described above is appropriate when the adiabatic condition $\hbar\omega_{0}\ll E_{F}$ is satisfied.
+[…]
+Figure 3: Critical energy scales and pairing strength.
+(a) Fermi energy $E_{F}$ of the 2DES, expressed in units of PDT single-mode energy $\hbar\omega_{0}$ for ferroelectric $\text{PbTiO}_{3}$, plotted as a function of Fermi wavenumber $k_{F}$ normalized to the ferroelectric thickness $L_{f}$.
+The long-wavelength regime described in Fig.
+2 is accessed in the limit of low density, shown as circled region.
+(b) Pairing strength parameter $\lambda-\mu^{*}$ calculated for a linear 2DES in the long-wavelength approximation as a function of balancing parameter $\zeta$ defined in Eq.
+(11), shown for two typical values of Fermi velocity $v_{F}$.
+Here, carrier density was fixed at $n=1\times 10^{15}/\text{m}^{2}$.
+Inset highlights where typical material systems (NC/DE) fall, according to the models used in this work.
+In frequency space, the negative permittivity associated with the PDT response [Eq.
+(5)] persists below the characteristic PDT oscillation frequency $\omega_{0}$ [39].
+Consequently, the Coulomb interaction mediated by the NC ferroelectric remains attractive for $\omega<\omega_{0}$.
+When the Fermi energy satisfies $E_{F}\gg\hbar\omega_{0}$, this separation of energy scales allows the retarded attractive interaction to partially overcome the instantaneous Coulomb repulsion, analogous to phonon-mediated pairing.
+Motivated by this observation, we next estimate the pairing strength within our model platform.
 Fig.
 3(a) shows the ratio of 2DES Fermi energy $E_{F}$ to PDT oscillation energy $\hbar\omega_{0}$ (for $4~\text{nm}$ thick $\text{PbTiO}_{3}$) as a function of carrier density.
-At low density, the adiabatic condition begins to be satisfied ($E_{F}\sim 10\hbar\omega_{0}$) only for linear 2DES with high $v_{F}$.
+At low density, this condition begins to be satisfied ($E_{F}\sim 10\hbar\omega_{0}$) only for linear 2DES with high $v_{F}$.
 At high density, both linear and parabolic 2DES comfortably satisfy this condition ($E_{F}\sim 100\hbar\omega_{0}$) for typical $v_{F},m^{*}$ values.
 The theory developed in this work is most applicable in the low-energy, long-wavelength regime.
-For this reason, we focus in the remainder of this section on the case of linear 2DES at low density, which according to Fig.
-3(a) could be treated as a low-energy system for large enough $v_{F}$ and according to Fig.
-2(b) could be modeled reasonably in the long-wavelength approximation (using Eq.
-(3)).
-Figure 3: Critical energy scales and pairing strength.
-(a) Fermi energy $E_{F}$ of the 2DES, expressed in units of PDT single-mode energy $\hbar\omega_{0}$ for ferroelectric $\text{PbTiO}_{3}$, plotted as a function of Fermi wavenumber $k_{F}$ normalized to the ferroelectric thickness $L_{f}=4~\text{nm}$.
-The long-wavelength regime described in Fig.
-2 is accessed in the limit of low density, shown as encircled region.
-(b) Pairing strength parameter $\lambda-\mu^{*}$ calculated for a linear 2DES in the long-wavelength approximation as a function of balancing parameter $\zeta$ defined in Eq.
-(13), shown for two different values of Fermi velocity $v_{F}$.
-Here, carrier density was fixed at $n=1\times 10^{15}/\text{m}^{2}$.
-Inset highlights where typical material systems (NC/DE) fall, according to the models used in this work.
+For this reason, we focus in the remainder of this section on the case of linear 2DES at low density, which we model in the long-wavelength approximation using Eq.
+(3).
 From the perspective of real systems, the high density regime and case of parabolic 2DES are just as likely to exhibit interesting physics, but careful treatment of $V_{\text{eff}}(q,\omega)$ in the short-wavelength, high-energy regime is required and beyond the scope of this work.
-Pairing strength: To quantify the strength of low-energy electron pairing in the 2DES, we use the parameters $\lambda$ and $\mu^{*}$, adapted from the theory of electron-phonon interaction and superconductivity [colemanbook].
+To quantify the strength of low-energy electron pairing in the 2DES, we use parameters $\lambda$ and $\mu^{*}$, adapted from the theory of electron-phonon interaction and superconductivity [8].
 To explain their definition in our context, we begin by decomposing the self-consistently screened interaction $V_{\text{eff}}$ in the following manner, based on the expected slowness of NC-mediated screening relative to direct screening by electrons ($\hbar\omega_{0}\ll E_{F}$):
@@
 We call the ‘slow’ term in Eq.
-(11) $\tilde{V}_{\text{env}}$ and calculate the following dimensionless constants [colemanbook]:
-$\begin{aligned} \lambda=-N_{0}\langle\tilde{V}_{\text{env}}\rangle=&-N_{0}\times\frac{\int_{0}^{2\pi}d\theta~\tilde{V}_{\text{env}}(q(\theta))}{2\pi},\\ \mu=N_{0}\langle V_{el}\rangle&=N_{0}\times\frac{\int_{0}^{2\pi}d\theta~V_{el}(q(\theta))}{2\pi},\\ \mu^{*}=&~\frac{\mu}{1+\mu\ln(\frac{E_{F}}{\hbar\omega_{0}})},\end{aligned}$ (12)
+(9) $\tilde{V}_{\text{env}}$ and calculate the following dimensionless constants [8]:
+$\lambda=-N_{0}\langle\tilde{V}_{\text{env}}\rangle\hskip 1.28032pt;\hskip 1.28032pt\mu=N_{0}\langle V_{el}\rangle\hskip 1.28032pt;\hskip 1.28032pt\mu^{*}=\frac{\mu}{1+\mu\ln(\frac{E_{F}}{\hbar\omega_{0}})},$ (10)
 where angle brackets denote averaging over Fermi surface (FS), $q(\theta)$ is appropriately defined for the 2D system, and $N_{0}$ is the density of states at $E_{F}$ per spin per unit area.
-The Coulomb pseudopotential $\mu^{*}$ [morelanderson] represents a renormalization of direct Coulomb repulsion ($\mu$) due to the large bandwidth for electron-electron scattering ($\sim E_{F}$) compared to environment-mediated screening ($\sim\hbar\omega_{0}$).
+The Coulomb pseudopotential $\mu^{*}$ [41] represents a renormalization of direct Coulomb repulsion ($\mu$) due to the large bandwidth for electronic screening ($\sim E_{F}$) compared to environment-mediated screening ($\sim\hbar\omega_{0}$).
 Carrier pairing at accessible temperatures is expected when $\lambda-\mu^{*}>0$ and sufficiently large ($\gtrsim 0.1$).
-In the long-wavelength regime, the problem can be fully specified by fixing values for the 2DES $n$ and $v_{F}$, and introducing a balancing parameter $\zeta$ defined as follows:
+In the long-wavelength limit, the problem can be fully specified by fixing values for the 2DES $n$ and $v_{F}$, and introducing a balancing parameter $\zeta$ defined as follows:
 $C_{nc}+C_{d}=-\zeta C_{q}.$ (11)
-To satisfy the long-wavelength stability condition in Eq. (5) we assume $\zeta>1$.
+To satisfy the thermodynamic stability condition in Eq.
+(4) we assume $\zeta>1$.
 Fig.
-3(b) shows calculated values of $\lambda-\mu^{*}$ for NC $\text{PbTiO}_{3}$ and linear 2DES ($n=1\times 10^{15}~/\text{m}^{2}$) in the long-wavelength regime.
+3(b) shows calculated values of $\lambda-\mu^{*}$ for NC $\text{PbTiO}_{3}$ and linear 2DES ($n=1\times 10^{15}~/\text{m}^{2}$) in the long-wavelength limit.
 We emphasize two points from this figure.
-First, the values of $\lambda-\mu^{*}\sim 0.2-0.3$ shown in the inset, calculated using the models described in previous section, are sufficiently large to motivate experimental efforts to observe pairing in such systems.
-For comparison, in monolayer graphene the value of $\lambda$ arising from other pairing mechanisms such as the electron-phonon interaction is many orders of magnitude smaller [einenkel2011graphene].
-Second, balancing the capacitances of NC, DE and 2DES offers a new route to engineer the value of $\lambda-\mu^{*}$.
-Particularly, as the balancing parameter approaches $\zeta=1$ from above, pairing can be pushed into stronger coupling regime.
-Design considerations: From the perspective of NC stabilization, linear and parabolic 2DES systems present qualitative differences.
-For a linear 2DES, $C_{q}\propto\sqrt{n}/v_{F}$, so the choice of carrier density directly affects the condition for NC stabilization (Eq.
-(5)).
-For a parabolic 2DES, $C_{q}$ is independent of carrier density, and scales with effective mass as $C_{q}\propto m^{*}$.
-This means that effective mass is the critical parameter and that platforms with a larger $m^{*}$ parabolic 2DES will require larger negative $\varepsilon^{nc}$ for NC to be stabilized.
-From the perspective of carrier pairing represented by $\lambda-\mu^{*}$, the functional dependence on $n$ and $L_{f}$ is more complicated due to the expressions for FS averaging and retardation effect ($\mu^{*}$).
-The essential route to maximizing interaction strength is to balance the positive and negative contributions in Eq.
-(2), and in the long-wavelength regime the problem simplifies considerably, leading to the points made in Fig.
-3(b).
-V Discussion
-In summary, we have presented a framework for Coulomb engineering using negative capacitance materials.
-Key results include: the idea of Coulomb Engineering using NC (Fig.
-1), a stability condition for NC in MF2IM configuration Eq.
-(5), categorization into regimes based on critical length scales (Fig.
-2) and an estimation of energy scales and interaction strength for models of material systems (Fig.
+First, estimates for the model platform introduced above yield $\lambda-\mu^{*}\gtrsim 0.1$ (inset).
+Although these estimates rely on simplifying assumptions, discussed further in the next section, they suggest that pairing strengths of experimentally relevant magnitude may be achievable.
+For comparison, in monolayer graphene the value of $\lambda$ arising from other pairing mechanisms such as the electron-phonon interaction is many orders of magnitude smaller [12].
+Second, in contrast to phonon-mediated pairing where the coupling strength is typically fixed by material properties, balancing the capacitances of NC, DE and 2DES offers a new route to engineer the value of $\lambda-\mu^{*}$.
+Particularly, as the balancing parameter approaches $\zeta=1$ from above, pairing can be pushed into strong coupling regimes [^1] We note that the assumptions of linear response and the existence of a small parameter $\lambda\hbar\omega_{0}/E_{F}$, which underlie the present pairing estimates, may break down as $\zeta\rightarrow 1$.
+This could ultimately limit the achievable pairing strength..
+Discussion.— In summary, we have introduced a new method for Coulomb engineering using negative-capacitance materials.
+We show that incorporating NC media enables the effective electron–electron interaction in two-dimensional systems to be tuned and even reversed in sign while maintaining thermodynamic stability (Fig.
+1).
+We additionally present a stability condition for NC in the MF2IM configuration [Eq.
+(4)], categorization of Coulomb engineering into regimes based on critical length scales(Fig.
+2), and estimates of interaction strengths for realistic material platforms (Fig.
 3).
-We now outline the limitations of our model and discuss directions for future work.
-First, the macroscopic Coulomb interaction $V_{\text{eff}}(\vec{q},\omega)$ is generally both wavevector and frequency dependent, and models of this function could be analyzed within a full Green’s function formalism for more quantitative analysis.
-Our model is best suited for the long-wavelength, static limit $V_{\text{eff}}(q\to 0,\omega\to 0)$, where we have additionally assumed that $V_{\text{eff}}$ is isotropic along the two in-plane directions of 2DES.
-In the context of engineered phase transitions, detailed knowledge of $V_{\text{eff}}(\vec{q},\omega)$ is required to predict which ordered ground state would be most stable.
-In the specific case of superconducting ground state, anisotropy of $V_{\text{eff}}$ in the plane of 2DES will affect whether the pairing is expected to be s-wave (as assumed here) or unconventional.
-Second, the model for coupling between 2DES and PDT (Kittel model) in Eqs.
-(6),(7),(8) could be improved in two steps: (i) modeling of optic branch dispersion $\omega_{+}(q)$ (see Appendix C) and coupling $\Omega(q)$ to extend the present result for $\varepsilon^{f}_{z}(q)$ to finite $q$, and (ii) calculation of contribution from all types and branches of excitations in the PDT system [sidorkin2006domain], amounting to complete knowledge of $\varepsilon^{f}_{z}(\vec{q})$ in terms of $\omega_{i}^{L(T)}(\vec{q})$ and $\Omega_{i}^{L(T)}(\vec{q})$ for each branch $i$ and type - longitudinal (L) or transverse (T).
-Third, the description of NC in $\text{PbTiO}_{3}$ based on the simple Kittel model may be inadequate when domain structures in experiments exhibit greater complexity (e.g., vortex textures [yadav2019]).
-Theoretical modeling of the collective excitations in such structures [yang2021condensation] is necessary to account for such details.
-Fourth, the models used for 2DES and domain-provided NC are essentially zero-temperature theories.
-This could be extended by (i) using the Lindhard function at finite temperature [giuliani2008quantum] for $\Pi_{el}$ in Eq.
-(2) and (ii) incorporating temperature-dependent domain configurations, which would effectively make the NC response function temperature-dependent [pavlenko2021temperature].
-Fifth, properties of the bare 2DES itself are expected to deviate significantly from the non-interacting description (used in this work) at low carrier densities.
-One well-known deviation is negative compressibility [eisenstein1992negative], in which case our assumption of a strictly positive $C_{q}$ from the non-interacting theory may not be a reliable approach.
-These corrections could be included in the present model by replacing $C_{q}$ with an expression that incorporates interaction effects [giuliani2008quantum].
-Finally, we note that the superconducting transition temperature $T_{c}\sim\exp(-1/(\lambda-\mu^{*}))$ so that small errors in estimating the quantity $\lambda-\mu^{*}$ can cause large errors in estimating $T_{c}$.
-This makes it important to design experiments that can measure small values of $\lambda-\mu^{*}$ directly.
-VI Acknowledgements
-SD would like to thank Sayeef Salahuddin for alerting him to transient NC and for helpful discussions regarding negative permittivity in ferroelectrics.
+We now outline the limitations of our analysis and discuss directions for future work.
+A more quantitative description of $V_{\mathrm{eff}}(\vec{q},\omega)$ requires extending the present treatment beyond the long-wavelength and low-frequency limits considered here.
+For the ferroelectric NC medium, this includes: (i) incorporating dispersion of the optic branch $\omega_{+}(q)$ and coupling $\Omega(q)$ to extend the present results for $\varepsilon_{z}^{f}(q=0)$ to finite $q$ (see Appendix C); (ii) including contributions from all excitation branches in the PDT system [55]; (iii) accounting for anisotropy introduced by domain structures, which may lead to angular dependence of $\epsilon_{z}^{f}(\vec{q})$ and influence pairing symmetry; (iv) extending the description of NC beyond the Kittel model [62], particularly given experimentally observed vortex-like structures [61]; and (v) incorporating the temperature dependence of the NC phase and its dielectric response [45].
+For the 2DES, further refinements include incorporating finite-thickness corrections through $q$-dependent form factors [1] and extending $\Pi_{\rm el}$ beyond the static, zero-temperature limit by using finite-temperature dynamical screening approaches [16].
+This extension incorporates the plasmon, which is important even in the low-frequency limit due to its gapless nature in two dimensions.
+Hybridization [23] of this mode with excitations of the NC medium can modify $V_{\mathrm{eff}}(\vec{q},\omega)$ and conclusions about the resulting ground states.
+Finally, the estimates of pairing strength using Eq.
+(10) can be improved by analyzing the full momentum- and frequency-dependent interaction $V(\vec{q},\omega)$ within a Green’s function framework [14, 12].
+When pair-breaking processes limit superconductivity, then the transition temperature scales as $T_{c}\sim\exp(-1/(\lambda-\mu^{*}))$, so that small errors in estimating the quantity $\lambda-\mu^{*}$ can cause large errors in estimating $T_{c}$.
+This makes it important to design experiments capable of measuring small values of $\lambda-\mu^{*}$ directly.
+Moreover, in two dimensions $T_{c}$ may instead be limited by phase fluctuations, highlighting the need to estimate the superfluid stiffness [50] in future work.
+Acknowledgements.—SD would like to thank Sayeef Salahuddin for alerting him to transient NC and for helpful discussions regarding negative permittivity in ferroelectrics.
 He would also like to thank Bhaskaran Muralidharan, Kerem Camsari and Zubin Jacob for their feedback on an earlier version of the manuscript.
@@
 This work was supported by the Purdue Research Foundation and the Blue Sky Research Program through the Purdue College of Engineering.
-Appendix A Anisotropic gate screening effect with Negative Capacitance
-We consider a two-dimensional electron system (2DES) encapsulated on one side by a dielectric medium with homogeneous positive static permittivity tensor, $\varepsilon^{d}=\mathrm{diag}(\varepsilon_{\perp}^{d},\varepsilon_{\perp}^{d},\varepsilon_{z}^{d})$, and on the other side by a negative capacitance (NC) medium described by $\varepsilon^{nc}=\mathrm{diag}(\varepsilon_{\perp}^{nc}>0,\varepsilon_{\perp}^{nc}>0,\varepsilon_{z}^{nc}<0)$, where $z$ denotes the stacking direction.
+Appendix A APPENDIX
+Appendix B A. Anisotropic gate screening effect with Negative Capacitance
+We consider a two-dimensional electron system (2DES) encapsulated on one side by a dielectric medium with homogeneous positive static permittivity tensor, $\varepsilon^{d}=\mathrm{diag}(\varepsilon_{\perp}^{d},\varepsilon_{\perp}^{d},\varepsilon_{z}^{d})$, and on the other side by a negative-capacitance (NC) medium described by $\varepsilon^{nc}=\mathrm{diag}(\varepsilon_{\perp}^{nc}>0,\varepsilon_{\perp}^{nc}>0,\varepsilon_{z}^{nc}<0)$, where $z$ denotes the stacking direction.
 From Poisson’s equation, the electrostatic potential $\varphi$ in the two encapsulating regions takes the form:
@@
 with anisotropy factors $\eta_{d}$ = $\sqrt{\varepsilon^{d}_{\perp}/\varepsilon^{d}_{z}}$ and $\eta_{nc}$ = $\sqrt{\varepsilon^{nc}_{\perp}/\varepsilon^{nc}_{z}}$.
-Metal gates are placed at distances $L_{d},L_{nc}$ above and below the $z=0$ plane where 2DES is located.
+Metal gates are placed at distances $L_{d},L_{nc}$ above and below the $z=0$ plane where the 2DES is located.
 Imposing the boundary condition $\varphi(q,L_{d/nc})$ = 0 at the gates and enforcing continuity of $\varphi$ across the 2DES plane, we obtain:
-$\begin{aligned} \varphi(q)\equiv\varphi(q,z=0)&=\alpha^{(nc)}(1-e^{-2\eta_{nc}qL_{nc}})\\ &=\beta^{(d)}(1-e^{-2\eta_{d}qL_{d}}).\end{aligned}$ (A2)
+$\begin{aligned} \varphi(q)\equiv\varphi(q,z=0)&=\alpha^{(nc)}(1-e^{-2\eta_{nc}qL_{nc}})\\ &=\beta^{(d)}(1-e^{-2\eta_{d}qL_{d}}).\end{aligned}$ (13)
 […]
-$\begin{aligned} &\alpha^{(nc)}\varepsilon_{z}^{nc}\partial_{z}(e^{\eta_{nc}qz}-e^{-2\eta_{nc}qL_{nc}}e^{-\eta_{nc}qz})\big|_{z=0}\\ &-\beta^{(d)}\varepsilon_{z}^{d}\partial_{z}(e^{-\eta_{d}qz}-e^{-2\eta_{d}qL_{d}}e^{\eta_{d}qz})\big|_{z=0}=\frac{e}{\varepsilon_{0}}.\end{aligned}$ (A3)
-Solving for $\alpha^{(nc)}$ and substituting in Eq. (A2), the potential in 2DES is:
+$\begin{aligned} \alpha^{(nc)}&\varepsilon_{z}^{nc}\partial_{z}(e^{\eta_{nc}qz}-e^{-2\eta_{nc}qL_{nc}}e^{-\eta_{nc}qz})\big|_{z=0}\\ -&\beta^{(d)}\varepsilon_{z}^{d}\partial_{z}(e^{-\eta_{d}qz}-e^{-2\eta_{d}qL_{d}}e^{\eta_{d}qz})\big|_{z=0}=e/\varepsilon_{0}.\end{aligned}$ (14)
+Solving for $\alpha^{(nc)}$ and substituting in Eq. (13), the potential in 2DES is:
 $\varphi(q)=\frac{e}{\varepsilon_{0}}\times\dfrac{1}{\varepsilon_{z}^{d}(q)\dfrac{\eta_{d}q}{\tanh(\eta_{d}qL_{d})}+\varepsilon_{z}^{nc}\dfrac{\eta_{nc}q}{\tanh(\eta_{nc}qL_{nc})}}.$ (15)
@@
 Finally, we note that for anisotropic NC media $\eta_{nc}$ is purely imaginary, but the static quantities $\varphi(q)$, $V_{\text{env}}(q)$ (and $V_{\text{eff}}(q)$ in main text) remain strictly real.
-Appendix B Stability of Negative Capacitance in MF2IM configuration
+Appendix C B. Stability of Negative Capacitance in MF2IM configuration
 We consider a system where the dielectric is held at voltage $V_{d}$, the NC medium at $V_{nc}$, and the 2DES at $V_{q}$. For small charge fluctuations, the free energy of the system is
@@
 Assuming $C_{nc}<0$ and $C_{d},C_{q}>0$, this condition simplifies to the stability condition quoted in the main text as Eq. (4).
-Appendix C Domain-provided Negative Capacitance model
-In a ferroelectric with periodic domain texture (PDT), the ‘restoring force’ characterizing the response of domain walls to an external electric field can be calculated within the Kittel model of alternating ‘hard’ domains.
-The equilibrium domain width in this model [catalan2012domain] is:
+An important consideration neglected in this work is the modification of $C_{q}$ (and $\Pi_{el}$ in main text) by ‘direct’ electron-electron interactions (the screening channel constituted by identical electrons within the 2DES).
+The strength of this interaction relative to kinetic energy is often expressed in terms of the dimensionless parameter $r_{s}$.
+For a 2DES with spin and valley degeneracies $g_{s}$ and $g_{v}$ and background dielectric constant $\kappa$, we have the following expressions for different cases of dispersion relation [9]:
+$r_{s}=\begin{cases}\displaystyle\frac{e^{2}}{4\pi\varepsilon_{0}\kappa\hbar v_{F}}\,\frac{\sqrt{g_{s}g_{v}}}{2},&\text{Linear},\\[6.0pt] \displaystyle\frac{m^{*}e^{2}}{8\pi\varepsilon_{0}\kappa\hbar^{2}}\,\frac{g_{s}g_{v}}{\sqrt{\pi n}},&\text{Parabolic}.\end{cases}$ (20)
+For a 2DES at large enough $r_{s}$, the modified $C_{q}$ could itself be negative due to exchange-correlation corrections [13, 16].
+In that case, Eqs.
+(17) and (18) remain valid, but Eq.
+(19) becomes only a necessary (not sufficient) condition, and Eq.
+(4) of the main text must be revised.
+A more detailed analysis indicates that, within the model used in this section, an MF2IM structure with both $C_{nc}<0$ and $C_{q}<0$ would actually be unstable, in the sense that no simple necessary and sufficient stability condition can be formulated.
+It is nevertheless conceivable that an appropriately designed stabilizing circuit could enable access to this regime, should it prove to be experimentally relevant.
+Appendix D C. Domain-provided Negative Capacitance model
+In a ferroelectric with periodic domain texture (PDT), the ‘restoring force’ characterizing the response of domain walls to an external electric field can be calculated within the Kittel model [33] of alternating ‘hard’ domains.
+The equilibrium domain width in this model [7] is:
 $d=\sqrt{3.53\sqrt{\frac{\varepsilon^{f}_{\perp}}{\varepsilon^{f}_{z,hf}}}\xi\delta L_{f}},$ (21)
-where $\xi=2\times(1+\varepsilon^{d}_{z}/\sqrt{\varepsilon^{f}_{\perp}\varepsilon^{f}_{z,hf}})$ is a parameter capturing electrostatic boundary conditions, $\delta\approx 1~\text{nm}$ is the domain wall thickness [luk2018prb], and $L_{f}$ is the ferroelectric thickness.
+where $\xi=2\times(1+\varepsilon^{d}_{z}/\sqrt{\varepsilon^{f}_{\perp}\varepsilon^{f}_{z,hf}})$ is a parameter capturing electrostatic boundary conditions, $\delta\approx 1~\text{nm}$ is the domain wall thickness [39], and $L_{f}$ is the ferroelectric thickness.
 The PDT system has a ‘stiffness’ because there is an energy cost from long-range Coulomb forces when this system is displaced from equilibrium.
-The oscillation frequency corresponding to this stiffness is calculated using Fourier analysis:
-$\begin{aligned} \omega_{\pm}^{2}&(q)=\frac{4P_{s}^{2}}{\pi\varepsilon_{0}\sqrt{\varepsilon^{f}_{\perp}\varepsilon^{f}_{z}}ML_{f}}\times\\ &\Bigg[\sum_{n=1}^{\infty}\textrm{ln}\bigg(1+\frac{\varepsilon^{f}_{\perp}}{\varepsilon^{f}_{z}}\frac{L_{f}^{2}}{d^{2}}\frac{1}{(2n-1)^{2}}\bigg)\big(1\pm\cos((2n-1)qd)\big)\\ &-\sum_{n=1}^{\infty}\textrm{ln}\bigg(1+\frac{\varepsilon^{f}_{\perp}}{\varepsilon^{f}_{z}}\frac{L_{f}^{2}}{d^{2}}\frac{1}{(2n)^{2}}\bigg)(1-\cos(2nqd))\Bigg],\end{aligned}$ (C2)
+The oscillation frequency corresponding to this stiffness is calculated in the Kittel model [55] using Fourier analysis:
+$\begin{aligned} \omega_{\pm}^{2}&(q)=\frac{4P_{s}^{2}}{\pi\varepsilon_{0}\sqrt{\varepsilon^{f}_{\perp}\varepsilon^{f}_{z}}ML_{f}}\times\\ &\Bigg[\sum_{n=1}^{\infty}\textrm{ln}\bigg(1+\frac{\varepsilon^{f}_{\perp}}{\varepsilon^{f}_{z}}\frac{L_{f}^{2}}{d^{2}}\frac{1}{(2n-1)^{2}}\bigg)\big(1\pm\cos((2n-1)qd)\big)\\ &-\sum_{n=1}^{\infty}\textrm{ln}\bigg(1+\frac{\varepsilon^{f}_{\perp}}{\varepsilon^{f}_{z}}\frac{L_{f}^{2}}{d^{2}}\frac{1}{(2n)^{2}}\bigg)(1-\cos(2nqd))\Bigg],\end{aligned}$ (22)
 where the two branches ($\pm$) of excitation for this system are analogous to the acoustic and optic branches in crystals with a two-atom basis.
-In this work we restrict attention to the ++ (‘optic’) branch in the long-wavelength limit, $q\to 0$, so $\omega_{0}$ in the main text is just $\omega_{+}(q=0)$.
+In this work we restrict attention to the ++ (‘optic’) branch in the long-wavelength limit, $q\to 0$.
+We calculate $\omega_{0}$ in the main text as $\omega_{+}(q=0)$ and assume electric dipole coupling of the form $P_{z}\times E_{z}$ [55, 39]:
+$\begin{aligned} \omega_{0}^{2}=\frac{8P_{s}^{2}}{\pi\varepsilon_{0}\sqrt{\varepsilon^{f}_{\perp}\varepsilon^{f}_{z,\text{hf}}}ML_{f}}&\mathrm{ln}\Bigg(\mathrm{cosh}\Bigg(\sqrt{\frac{\varepsilon^{f}_{\perp}}{\varepsilon^{f}_{z,\text{hf}}}}\frac{\pi}{2}\frac{L_{f}}{d}\Bigg)\Bigg)\\ \Omega^{2}=&\dfrac{4P_{s}^{2}}{\varepsilon_{0}\varepsilon^{f}_{z,\text{hf}}Md}.\end{aligned}$ (23)
+Parameters for ferroelectric $\text{PbTiO}_{3}$ are chosen as follows [39] : $P_{s}=0.65~\text{Cm}^{-2}$ is the spontaneous polarization, and $M$ is the domain wall mass per unit area estimated using interpolation formula $M=1.3\sqrt{L_{f}[\text{nm}]}\times 10^{-9}~\text{kgm}^{-2}$ [39, 64].
+$\varepsilon^{f}_{z,hf}=100$ is the high-frequency dielectric constant capturing the background polarizability of $\text{PbTiO}_{3}$ arising from other polarization mechanisms in the material, and $\varepsilon^{f}_{\perp}=30$.
 Within this model, the calculated negative permittivity $\varepsilon^{f}_{z}$ has a value of about 60-60 for $L_{f}=4~\text{nm}$.